A Complete Bibliography of Journal of the Royal Statistical Society, Series D (The Statistician):

1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.01

Title word cross-reference

(p, q) [85]. ^{(}R) [721]. A [655, 181]. α [134]. β [134]. χ^{2} [454]. D [181]. g [655]. G^{2} [454]. H [63]. K [375, 718]. N [134]. p_{1} [134]. p_{2} [134]. S [381]. t [532, 123]. T^{2} [600]. W′ [205].

-Class [718]. -Function [63]. -Optimal [181]. -Sample [375]. -Shaped [381]. -Test [123]. -Tests [532].

1 [449]. 1900 [558]. 1970 [224]. 1979 [83].

29 [434].

3 [721, 484].

40 [448]. 43 [433]. 45 [434].

53 [472]. 5497 [523]. 57 [510].

8 [731].

Abstract [176]. Abuse [177, 411]. Acceptability [594]. Acceptance [589, 635, 814, 176]. Accident [184, 732, 171]. Accidents [614, 405]. Accounting [766]. Accuracy [350]. Achen [434]. Active [723]. Addendum [695, 463]. Additive [377]. Address [608]. Adjust [88]. Adler [790, 790]. Advanced [710, 347, 577, 449]. Advances [243, 582]. Advertising [419, 416]. Africa [226, 645, 216]. African [122]. after [613]. Age [411]. Agricultural [297]. Ahlswede [704]. Aickin [442, 621]. Aid [283, 130, 595, 460]. Aided [197]. Aids [106, 602]. Aitchison [573]. Akaike [698]. Alan [622, 755, 394]. Albert [93, 349]. Aldrich [434]. Alexander [124]. Alfred [759, 714]. Algebra [143, 812]. Algorithm [690, 402, 514]. Algorithms [270, 474]. Ali [757]. Alice [342]. Allan [735]. Allen [311]. Allocation [507]. Allocations [64]. Alternative [494, 691, 269, 607, 53, 209, 523, 646, 153, 609, 272]. Alternatives [688, 532]. Alvey [728]. Ambartzumian [239]. Amendola [759]. American [813]. Analyse [501]. Analyses [169, 682, 747]. Analysing [331, 51]. Analysis [365, 207, 633, 520, 327, 43, 195, 15, 261, 442, 781, 467, 27, 173, 619, 38, 470, 573, 58, 204, 563, 733, 753, 526, 42, 481, 72, 234, 120, 238, 561, 230, 383, 472, 754, 460, 209, 665, 356, 147, 332, 182, 77, 221, 562, 540, 757, 453, 79, 789, 555, 13, 538, 409, 206, 705, 135, 521, 411, 170, 418, 296, 667, 12, 510, 186, 528, 707, 172, 321, 387, 380, 163, 190, 570, 568, 124, 480, 156, 565, 474, 580, 697, 96, 295, 145, 735, 500, 569, 676]. Analytic [430, 454]. Analytical [508]. Andersen [350]. Anderson [81, 705, 666, 813]. Andrew [781]. Anemia [402]. Angular [58]. Aniello [759]. Animal [724]. Ann [430]. Annick [756]. Announcement [395]. ANOVA [478]. Answers [785]. AOQL [367]. AOQL-System [367]. AoV [89]. Applicable [812]. Application [192, 537, 120, 460, 801, 206, 255, 380]. Applications [589, 635, 814, 239, 519, 539, 448, 46, 559, 472, 77, 799, 157, 63, 434, 582, 166, 433, 510, 429, 348, 796, 163, 31, 389, 702, 629, 498, 220, 807]. Applied [184, 365, 478, 10, 74, 200, 67, 524, 430, 430, 388, 244, 70, 615, 102]. Appraisal [687]. Approach [522, 692, 651, 400, 175, 193, 187, 20, 401, 392, 546, 636, 476, 592, 172, 430, 575, 561, 810, 359]. Approaches [768, 209, 763, 646, 611, 258, 194]. Approximate [188, 461]. Approximating [191, 132, 606]. Approximation [452]. Approximations [414, 688, 800, 75]. Arc [378, 75]. Archive [497]. Area [752, 469, 90, 321]. Ariane [81]. Ariav [736]. ARMA [85]. Armitage [79]. Armstrong [100]. Arndt [241]. Arnold [556]. Arthanari [146]. Arthur [726]. Artificial [672]. Arya [216]. Ash [643]. Asia [226]. Asmussen [615]. Aspects [333, 334, 288, 105, 317, 654, 279]. Assays [364]. Assess [86]. Assessing [227]. Assessment [183, 180, 523, 342]. Assessments [610]. Assessors [331]. Assignments [664]. Associated [778, 493, 643, 738]. Association [550, 349]. Assumption [723]. Assurance [592, 601, 171]. Asymptotic [214, 743, 727]. Attacks [332]. Attenuation [424]. Audience [403]. Auditing [175, 203, 596, 797]. Auquier [81]. Autocorrelation [662]. Automatic [647]. Automation [765]. Automobile [167]. Autoregression [695, 649]. Autoregressive [329, 647]. Availability [196]. Aven [665]. Avon [264]. Awareness [419]. Azlarov [738].

B [142, 671, 768, 725, 712, 733, 664, 734, 18, 386, 565, 383, 696, 727, 217, 243, 714, 385, 64, 145, 51, 224, 226]. Back [9, 24, 28, 41, 47, 62, 78, 104, 116, 127, 148, 159, 201, 211, 249, 265, 276, 290, 310, 325, 339, 360, 371, 396, 420, 450, 464, 485, 512, 541, 585, 604, 637, 652, 678, 716, 740, 774, 792, 816, 334]. Bailey [139, 277]. Ball [100]. Bancroft [766]. Band [424]. Bandit [507]. Barbara [435]. Barnett [441, 99]. Barrie [557]. Bartholomew [7]. Basawa [217]. Based [610, 268, 439, 476, 258, 684, 301, 796].

Bases [759]. Basic [666, 725, 597, 8, 668, 242, 709, 345, 749, 498, 340]. Basics [478]. Basil [504]. Basset [768]. Bassett [671]. Basu [571]. Batch [176]. Bateson [390]. Bauerle [788]. Bayes [336, 203, 724, 524, 181]. Bayesian [750, 184, 522, 692, 651, 175, 338, 196, 659, 174, 520, 195, 192, 187, 494, 516, 169, 539, 518, 535, 523, 161, 188, 534, 536, 182, 540, 532, 199, 407, 197, 166, 521, 210, 418, 399, 513, 194, 186, 406, 528, 172, 797, 380, 163, 433]. Bayesians [189]. BBC [668]. Be [315, 683]. Beaumont [564]. Beck [582]. Becker [702]. Before [558]. Behavioral [126]. Behaviour [19]. Belief [526]. Benefit [206]. Benefits [630]. Benjamin [805]. Berg [445]. Bergum [807]. Berkshire [90]. Bernard [805]. Berry [507, 582, 731]. Bert [507]. Bertil [620]. Best [786, 503]. Beta [183, 194]. Beta-Thalassaemia [194]. Better [630]. Betting [208]. Between [327, 454, 136, 60, 535, 366, 272, 194]. Beyond [478]. Biased [443]. BIASs [200]. Bibby [245]. Bifurcation [157]. Bifurcation-Applications [157]. Billingsley [471, 625]. Binary [179]. Binomial [723, 358, 461]. Bioequivalence [776, 611]. Biologists [804]. Biology [19, 139]. Biomathematics [277]. Biostatistics [670]. Bivariate [286, 374, 468]. Blackspot [184]. Blake [67]. Block [144, 789, 181]. Blood [402]. Boiler [261]. Book [140, 142, 567, 350, 710, 430, 504, 633, 784, 438, 772, 560, 627, 699, 785, 239, 21, 384, 432, 761, 144, 66, 146, 155, 499, 704, 216, 500, 471, 388, 441, 577, 475, 578, 730, 767, 769, 806, 812, 783, 297, 766, 671, 241, 768, 788, 666, 442, 617, 781, 220, 341, 725, 576, 482, 703, 437, 572, 674, 712, 706, 235, 242, 511, 17, 570, 575, 31, 218, 447, 558, 619, 709, 281, 298, 573, 223, 291, 292, 501, 568, 618, 345, 478, 563, 708, 733, 753, 732, 664, 668, 219, 807, 20, 81, 97]. Book [98, 124, 125, 141, 240, 729, 698, 16, 584, 622, 755, 244, 579, 481, 70, 734, 786, 19, 138, 448, 507, 669, 673, 143, 443, 446, 392, 480, 68, 101, 726, 749, 809, 389, 18, 346, 677, 630, 559, 69, 156, 238, 386, 439, 440, 505, 561, 565, 615, 620, 773, 623, 483, 632, 279, 503, 506, 383, 474, 581, 472, 484, 731, 754, 805, 813, 476, 580, 473, 665, 675, 756, 750, 95, 566, 696, 727, 752, 771, 711, 278, 147, 763, 621, 810, 765, 634, 344, 435, 102, 787, 390, 7, 67, 139, 217, 221, 243, 391]. Book [479, 562, 342, 670, 556, 702, 757, 701, 343, 751, 557, 629, 22, 157, 569, 697, 789, 477, 626, 431, 790, 225, 277, 63, 82, 300, 737, 498, 502, 625, 714, 762, 770, 385, 393, 434, 509, 764, 728, 280, 508, 582, 736, 705, 759, 758, 282, 299, 808, 359, 64, 65, 738, 444, 436, 449, 574, 301, 382, 433, 616, 760, 782, 99, 294, 296, 445, 624, 667, 676, 236, 237, 631, 811, 510, 564, 628, 713, 222, 245, 791, 96, 295, 571, 583, 707, 394, 804, 293, 349, 497, 8, 340, 83, 126, 246, 145, 51]. Book [672, 100, 387, 348, 347, 735, 700, 224, 226]. Bootstrap [546, 636]. Bootstrapping [776]. Bortkewitsch [660]. Boruch [630]. Bound [42]. Boundary [553]. Bounds [529]. Bowen [145]. Bowhead [268]. Box [616, 499, 69, 536]. Boyle [482]. Brake [83]. Branch [42]. Bray [562]. Breast [107]. Bremner [671]. Brewer [294]. Brian [578, 435]. Bridge [327]. Brillinger [475]. Britain [22, 83]. British [271]. Brittany [22]. Broad [424]. Brownian [506]. Bruce [502]. Brunberg [446]. BS [523]. BTO [509]. Build [344]. Building [616]. Bullen [773]. Bulletin [200]. Bunn [431]. Burke [560, 560]. Bus [86]. Business [66, 766, 242, 55, 345, 3, 46, 632, 293, 642, 645, 726, 702]. Butler [786, 786].

C [633, 438, 384, 216, 297, 768, 341, 725, 511, 753, 97, 622, 156, 386, 561, 675, 752, 102, 343, 697, 449, 628, 295, 707, 804, 672]. Calculating [522, 692, 651]. Calculation [1, 92, 304, 49, 48]. Calculus [577, 389]. Campbell [804]. Can [315, 683]. Cancer [107, 366]. Capture [659]. Capture-Recapture [659]. Car [405]. Carbone [350]. Cardiovascular [720]. Care [101]. Carey [584]. Carl [483]. Carlo [566, 718, 533, 61, 186, 406, 528, 190]. Carriers [194]. Carver [66]. Case [414, 430, 152, 657, 304, 460, 258, 722, 413]. Case-Control [413]. Cases [454, 645]. Catalogue [497]. Catalysis [768]. Catastrophe [157]. Catch [411]. Catch-at-Age [411]. Catcher [584]. Catchment [1, 92, 112, 49, 48]. Categorical [802]. Categories [331]. Caulcutt [240]. Causal [583]. Caused [86]. Caution [267]. Cautionary [26]. CBE [311]. Celia [449]. Cell [402, 717, 798]. Cells [195]. Cellular [19]. Census [34, 37, 813]. Censuses [55, 122]. Ceramic [687]. Chaghaghi [444]. Chained [272]. Challenges [783]. Change [329, 458, 192, 425, 469]. Change-Over [458]. Changes [54, 748, 428]. Changing [412]. Chapman [504]. Characteristics [260]. Characterization [73, 789, 738]. Characterizations [65]. Chart [128, 799]. Charts [50, 543, 328]. Chatfield [124, 761]. Chatterjee [757]. Checks [380]. Chemicals [763]. Chemistry [768, 168]. Chhikara [791]. Chi [229, 606, 468, 452, 75]. Chi-Square [229, 606, 468]. Chi-Squared [452, 75]. Child [177]. Children [316]. Chistyakov [64]. Choice [551, 553, 690, 199, 407, 399, 284]. Choosing [414, 272]. Chow [343]. Chris [20, 344]. Christensen [785]. Christopher [761, 124, 434]. Cigarette [366]. Claims [171]. Clark [101]. Clarke [8, 340]. Class [718, 735]. Classical [401]. Classification [517, 167, 334, 179]. Classroom [457]. Clay [204]. Clifford [736]. Clinical [525, 108, 180, 563, 109, 168, 105, 445, 106, 596, 359]. Closed [724, 461]. Cluster [552, 77].

Cluster-Analysis [77]. Clustering [218]. Coal [643]. Coalworkers [332]. Cochran [769, 296]. Code [608]. Coefficient [491, 684, 210]. Coffman [665]. Cognition [579]. Cohen [509]. Coherent [527]. Collins [124]. Column [424, 779]. Combinatorial [239]. Combining [354]. Comments [286, 49, 48]. Common [131]. Commonly [182]. Communication [598]. Communities [19]. Companies [722]. Company [525, 171]. Comparability [324]. Comparative [113, 81, 30, 126, 784, 431]. Comparing [378, 686, 658, 75, 134]. Comparison [207, 162, 270, 535, 695, 649, 452, 495, 12, 194, 798]. Comparisons [806, 755]. Compensating [268]. Competing [410]. Complex [785, 568, 808]. Complicated [332]. Component [365, 580, 12]. Components [88, 352, 303, 289, 488]. Compositional [573]. Compromise [712]. Computation [168, 514, 758]. Computational [388, 193, 561]. Computed [406]. Computer [44, 219, 665, 702, 197, 721, 100, 602, 796]. Computer-Aided [197]. Computer-Based [796]. Computing [746, 3, 758, 170, 528, 181, 340]. Conall [482]. Concepts [666, 725]. Condorcet [777]. Conference [338]. Confidence [686, 691, 546, 636, 524]. Confirmatory [285]. Conflict [155]. Conjectures [15]. Conjunction [689, 515]. Conover [392]. Considerations [610]. Constrained [356]. Construction [390]. Consumer [594, 136, 272, 803]. Consumption [262]. Contamination [168]. Content [643]. Contingency [550, 230, 459, 377]. Continuity [378, 75]. Continuity-Corrected [378, 75]. Contribution [259]. Control [686, 597, 367, 723, 600, 799, 119, 685, 413, 663, 602, 328, 384, 236, 237, 571]. Controlled [108, 133]. Controlling [314].

Cooke [340, 8]. Cooper [430]. Copper [337]. Corrected [378, 75]. Correction [50]. Corrections [651]. Correlated [656, 135]. Correlation [366, 684]. Corrigendum [814, 739, 695]. Corsi [343]. Costs [630, 405]. Council [264]. Counted [799]. Counterexamples [729]. Counting [238]. Countries [467, 55, 46, 304]. County [264]. Course [141, 8, 388]. Covariance [209, 356, 13, 369]. Cox [628]. Coxon [219, 278]. Craven [340]. Crime [38]. Criminology [429]. Criteria [272]. Criterion [698, 131]. Critical [319, 84]. Critique [123]. Croft [244]. Crops [784]. Cross [111, 731, 426]. Cross-Cultural [731]. Cross-Over [111]. Crossing [612]. Crossover [117]. Crouzeix [501]. Crystal [100]. Cultural [731]. Cumulative [606, 688, 799]. Current [260]. Curve [185]. Curves [659, 231, 381, 540]. CUSUM [799, 50, 695, 649, 543]. Cusums [363]. Cuthbert [707]. Cyclic [626].

D [142, 438, 66, 475, 578, 241, 788, 220, 235, 31, 733, 786, 143, 773, 383, 473, 696, 771, 147, 621, 102, 7, 702, 557, 789, 431, 625, 385, 434, 582, 510, 8, 340, 145]. Damned [245]. Daniel [707]. Data [530, 365, 408, 438, 353, 66, 227, 781, 341, 218, 58, 668, 42, 303, 72, 331, 156, 565, 474, 179, 472, 754, 460, 381, 221, 426, 799, 802, 682, 747, 643, 79, 697, 45, 409, 166, 705, 759, 411, 382, 295, 707, 172, 497, 145, 387, 602, 410, 743, 380, 441, 442, 573, 733, 673, 147, 751, 569, 301, 99, 811, 51, 700, 390]. Database [736]. Datasets [423]. Dates [338]. David [430, 619, 81, 244, 579, 446, 749, 782, 445, 510, 295, 806]. Davies [278, 219]. Davis [583]. Davison [346]. Debate [319]. Decennial [34]. Decision [322, 323, 461, 106, 672, 712, 634, 701]. Decision-Making [106]. Decisions [525, 46, 293]. Defence [589, 635, 814]. Deficiency [402]. Degradation [337].

Degree [454, 74]. Degroot [567]. Demand [657, 251, 252, 427, 255]. Deming [599]. Demystifying [68]. Dennis [93]. Densities [514]. Density [95, 565]. Dependence [60]. Dependences [301]. Dependent [780, 682]. Depletion [682, 747]. Deprivation [694]. Derived [334]. Describe [128]. Describing [182]. Description [353]. Design [430, 329, 152, 193, 703, 494, 563, 275, 305, 269, 480, 33, 221, 105, 317, 368, 328, 633, 707]. Designed [467]. Designing [525]. Designs [458, 779, 630, 288, 181, 144, 789, 626]. Detecting [417, 60]. Detection [192, 107, 233, 756]. Determining [796]. Developing [467, 55, 46, 304]. Development [19, 253, 240]. Developments [2, 356, 771]. Diaconis [459]. Diagnostic [113, 800]. Diagnostics [744, 407]. Diary [338]. Dichotomous [551]. Dieter [811]. Differ [331]. Difference [454, 691, 492, 546, 636]. Differenced [643]. Different [270]. Differential [706, 320, 681]. Différentielles [501]. Difficult [186]. Difficulty [689, 515]. Digits [45]. Dilemma [46]. Dillon [96]. Directions [736]. Directory [622]. Disciplines [63]. Discount [650]. Discovery [456, 797]. Discrete [442, 84, 653, 468, 96]. Discretised [743]. Discriminant [332, 96]. Discrimination [621]. Discussion [84]. Diseases [314]. Dispersion [170]. Display [550, 73, 393]. Distance [535]. Distribution [530, 710, 195, 183, 553, 606, 688, 355, 763, 402, 358, 461, 375, 210, 748, 308, 91, 142, 386, 738, 791]. Distributions [191, 493, 270, 470, 84, 653, 468, 168, 45, 170, 406, 528, 190]. Disturbances [718]. Division [39]. DLM [538]. Doane [782]. Dodge [146, 569]. Domain [475, 437]. Domains [33]. Donald [507, 483]. Doses [318]. Douglas [384]. Down [424]. Draper [616]. Drink [83]. Drinking [71]. Drug [364, 681]. Drugs [763]. DTI [722]. Duckstein [674]. Dummies [640, 453].

Dummy [551]. Duration [410]. Durbin [121]. Duties [14]. Dyke [784]. Dynamic [537, 188, 236, 237]. Dynamics [416, 737].

Early [107]. Econometric [447, 190, 31]. Econometrics [163, 220, 143]. Economic [216, 663]. Economically [256]. Economics [702, 242, 345, 632]. Edinburgh [317]. Edition [772]. Editor [80, 154, 115, 94, 23, 137, 463, 215]. Editorial [362, 398]. Education [517, 29]. Educational [699, 166, 51]. Edward [810, 393]. Edwin [447, 498]. Effect [306, 85, 59]. Effective [120]. Effectively [14]. Effects [128, 329, 780, 6, 681, 308]. Efficiency [152]. Efficient [168]. Efron [459]. Egon [436]. Ehrenberg [341]. Elandt [156]. Elandt-Johnson [156]. Election [271]. Electrical [431]. Electricity [252, 255]. Electrodiagnosis [113]. Elements [625, 714, 758]. Elicitation [183]. Eliot [810]. Elliott [622, 389, 449]. Elwyn [223]. Empirical [336, 657, 301, 524, 559, 616]. Empirically [334]. Employment [654]. Endrenyi [221]. Energy [259, 262, 250, 257, 460]. Engineering [577, 457, 674]. Enhancement [374]. Enhancements [130]. Enough [588]. Enquiry [436]. Enuresis [79]. Environmental [404, 86, 316, 342]. Enzyme [221]. Epidemiological [318, 316]. Epidemiology [342, 413]. Episodes [379, 748]. Epstein [243]. Equality [654]. Equations [808, 501]. Equilibrium [505]. Equivalence [723]. Eradicating [314]. Errata [635, 636]. Erratum [651]. Error [617, 643, 36]. Errors [128, 329, 613, 656, 44, 529, 382, 481, 246]. ESRC [722, 497]. ESRC/DTI [722]. Essential [473]. Establishing [775]. Estimates [85, 136, 555]. Estimating [455, 717, 377, 256, 112]. Estimation [650, 336, 662, 353, 547, 544, 647, 270, 229, 85, 274, 412, 565, 653, 33, 356, 203, 54, 795, 488, 724, 210, 661, 443, 95, 301]. Estimators [693, 739, 718, 801]. Ethics [484]. Ethnographic [754]. Ethnography [754]. European [208, 224]. Eurostat [772]. Evaluating [335, 205, 101, 343]. Evaluation [34, 318, 162, 337, 37, 260, 206, 368, 575]. Evans [68, 347]. Everitt [142]. Evidence [527]. Ewen [101]. Exact [546, 636]. Examination [88, 297, 352]. Examine [723]. Examining [10]. Example [548, 190]. Examples [656, 539, 53, 489, 123, 376]. Exchange [538]. Exiting [647]. Expectation [15, 402]. Expectation-Maximisation [402]. Expected [164, 544, 693, 739]. Expenditure [136]. Experience [720, 29, 457]. Experiment [149, 164, 297]. Experimental [152, 275]. Experimentation [133]. Experiments [365, 784, 193, 178, 569, 13, 633, 281, 298, 563, 507, 221, 383]. Expert [305, 344, 527, 672]. Explained [666]. Exploratory [230, 697, 295]. Exploring [423, 303, 438, 782, 445]. Exponential [496, 738, 176]. Exposed [318, 315]. Exposure [336]. Expressions [461]. Extended [176]. Extension [721]. External [772]. Extreme [530, 778].

F [438, 781, 712, 235, 345, 807, 97, 480, 630, 476, 435, 7, 67, 391, 670, 762, 705, 64, 449, 145, 672]. Fabian [767]. Factor [27]. Factorial [13]. Factors [720, 131]. Failure [598, 147]. Failures [182]. Farm [614]. Farmer [431]. Farmers [456]. FBA [311, 247]. Feasibility [107]. Feiring [502]. Fibonacci [807]. Field [784, 297, 731]. Fields [790]. Fieldwork [484, 754]. Fienberg [567]. Figures [504]. Fildes [350]. File [608]. Final [45]. Financial [581]. Finding [134]. Finite [369, 142, 474]. Firm [470]. Firm-Size [470]. First [329, 388, 410]. First-Passage-Time [410]. Fish [411]. Fisher [499, 69]. Fit [227, 425, 355, 468, 684]. Fitting [553, 381, 690]. Fixed [552]. Fleiss [563]. Flexible [521]. Floor [597]. Flow [506]. Flows [408]. Folks [791]. Following [329]. Football [208]. Forbes [101, 7]. Forecasters [162]. Forecasting [648, 350, 271, 641, 647, 518, 749, 251, 646, 252, 258, 255, 171, 100, 645, 726, 431]. Forecasts [254]. Forensic [689, 516, 534]. Form [461]. Formulae [214, 272]. Formulation [548]. Forrest [629, 434]. Foundations [754, 669]. Four [553]. Fowler [509]. Fox [348]. Frame [608, 720]. Frames [122, 722]. Framework [599]. Francia [205]. Francis [770]. Francois [444]. Frank [508]. Franken [241]. Frederic [500]. Frederick [296, 295]. Free [747]. Freed [497]. Freedom [454]. Freeman [575]. French [634]. Friends [707]. Fristedt [507]. Front [4, 25, 32, 35, 52, 57, 76, 87, 110, 118, 129, 160, 202, 213, 232, 266, 273, 302, 307, 313, 326, 351, 361, 372, 397, 421, 451, 465, 487, 542, 586, 587, 605, 638, 679, 680, 741, 742, 793, 794]. Fruit [365]. Fuel [264]. Fuller [617]. Function [612, 63]. Functional [274]. Functions [540, 377, 440, 476]. Fundamentally [533]. Fundamentals [483]. Further [234, 356, 79]. Future [259, 644, 260].

G [784, 216, 769, 766, 666, 220, 242, 709, 292, 478, 807, 698, 786, 143, 726, 18, 559, 754, 473, 665, 391, 557, 157, 697, 737, 65, 436, 574, 616, 296, 667, 676, 564, 713, 8, 340, 83, 672]. Gabor [764]. Gad [736]. Gagné [669]. Galwey [728]. Game [155]. Gamma [493]. Gandy [48]. Garstang [664]. Gary [734, 760]. Gas [657, 251]. Gatenby [668]. Gates [768]. Gaussian [791, 375]. GDP [136]. Gegerenzer [579]. Gene [556]. Genealogical [189]. General [271, 720, 109, 611, 748]. Generalised [723, 660]. Generating [440]. Genstat [623, 728]. Geoffrey [783]. Geographical [408]. Geometric [12]. Geometry [239]. Gerd [579]. Gilber [755]. Giles [619]. Giri [633]. Glazebrook [557]. GLIM [721, 711]. GLIM3 [5]. Godfrey [574]. Goldstein [96, 699]. Good [554]. Goodness [227, 425, 355, 468, 684]. Goodness-of-Fit [468, 684]. Graham [812]. Graphic [97, 73]. Graphical [550, 231, 84, 120, 536, 513, 697]. Graphics [721]. Gray [622]. Great [83]. Greater [683]. Gregory [343]. Gross [483]. Group [686, 748]. Grouping [126]. Groups [324, 686]. Growth [659, 231, 428]. Gruber [443]. Guba [436]. Gudmund [433]. Guide [219, 509, 624, 510, 761]. Gupta [789]. Guttman [131]. Gwyn [347].

H [567, 560, 699, 806, 575, 281, 298, 708, 753, 141, 138, 440, 561, 621, 562, 697, 225, 434, 676, 340, 126, 145, 672]. Hadi [757]. Hair [705]. Hamakers [688]. Hamer [629]. Hampel [476]. Hand [142]. Handbook [138, 628, 435, 475, 437, 97, 386, 479]. Handcrafting [572]. Handling [668, 802, 186]. Hanif [294]. Hanke [726]. Hannan [437, 767]. Hare [383]. Harnett [702]. Harold [280]. Harris [430, 483]. Harrison [506]. Hassett [730]. Hauck [81]. Hazard [612]. Hazards [800]. Health [322, 323, 319, 1, 92, 333, 320, 335, 770, 49, 48, 321]. Healy [711]. Henry [622]. Herbert [81, 382]. Heritability [555, 491]. Heteroscedasticity [613]. Hettmansperger [439]. Heyer [281, 298]. Hibon [350]. Hierarchical [519, 185, 528]. High [318]. Hills [79]. Historical [723, 224, 226]. History [558, 629, 813]. Hoaglin [438, 295]. Hoel [242]. Holder [477]. Hollinger [447]. Holmes [503]. Holt [641]. Honsberger [98]. Horadam [807]. Hormone [364]. Horse [660]. Horse-Kicks [660]. Hospital [112]. Housing [37, 262]. Howard [575]. Howson [783]. Huntsberger [625]. Hutchinson [732]. Hydrocarbon [256]. Hydrology [199, 734, 477].

Ian [577, 618, 734, 68]. IBM [782]. IDA [548]. Ideas [394]. Identifiability [531]. Identification [647, 231]. Identifying [84]. II [19, 754]. III [810]. Illness [379, 748]. Illustration [168]. Illustrative [376, 190]. Iman [392]. Impact [769]. Impacts [86]. Importance [514]. Improving [71, 677]. Imputation [696]. Imputing [545]. Incomplete [144]. Independent [378, 354]. Index [2, 306, 272]. Indices [772, 304, 214, 272, 803]. Indirect [151]. Industrial [128, 174, 304, 387]. Industry [373, 598, 3, 383, 590]. Industry-Some [373]. Inequalities [773]. Infectious [314]. Inference [560, 525, 516, 439, 217, 643, 433, 797, 675, 727, 625, 222]. Inferences [406]. Influence [549, 476]. Information [40, 103, 158, 248, 312, 370, 430, 486, 603, 715, 815, 323, 164, 698, 150, 416, 393, 673]. Informative [655, 173]. Initial [85]. Innovation [456]. Inspection [128]. Inspections [795]. Instructional [669]. Insurance [167]. Integral [239]. Integration [186, 406, 528, 225]. Integrative [430]. Intelligence [316, 672]. Inter [523]. Inter- [523]. Interchangeability [607]. Interfaces [788]. Interference [280]. Internal [336, 367]. International [752, 538, 226]. Interpretation [274, 383, 130]. Interpreting [434]. Intervals [686, 691, 546, 636, 685, 524]. Interviewing [754]. Intra [523, 491]. Intra-Laboratory [523]. Intra-Sire [491]. Intriligator [31]. Introducing [223]. Introduction [668, 634, 67, 713, 145, 781, 711, 502, 728, 384, 66, 767, 20, 124, 675, 670, 280]. Introductory [341, 510, 628, 730, 17, 762]. Intuitive [579]. Invariance [222]. Inventory [187]. Inverse [606, 688, 375, 791]. Inverse-Gaussian [375]. Investigation [53, 234]. Iraq [304].

Iron [402]. Irradiation [788]. Irvine [68]. Ishiguro [698]. Issues [641]. Italy [335]. Item [802]. Items [187]. IUCDs [337]. Iverson [676, 433].

J

[142, 567, 350, 710, 438, 21, 155, 730, 767, 769, 671, 768, 437, 674, 235, 242, 511, 573, 345, 563, 733, 668, 124, 141, 755, 579, 734, 138, 443, 392, 101, 389, 559, 386, 623, 279, 506, 383, 731, 813, 476, 675, 756, 752, 711, 147, 7, 139, 243, 562, 670, 626, 277, 385, 509, 764, 280, 705, 282, 299, 359, 574, 791, 126, 100, 387]. J.-M [768]. Jacob [382]. Jacobsen [238]. Jaech [481]. Jaeger [432]. James [95, 736, 583]. Jeff [68]. Jenkins [536]. Jessen [242]. Jill [472]. Joan [499, 69]. Job [198]. John [447, 481, 68, 726, 731, 434, 245, 295, 348, 144, 626]. Johnson [749, 156, 627, 156, 30]. Joiner [435]. Jolliffe [671, 480, 580]. Jones [671, 155, 508]. Jordan [729]. Joseph [446]. Jumps [188]. Jure [218]. Jury [777].

Kadane [567]. Kaiser [131]. Kalbfleisch [147]. Kapadia [666]. Kathleen [21]. Kaye [621]. Kazmier [345]. Keep [591]. Keith [388]. Keller [759]. Kendall [247, 710]. Kenwyn [445]. Kevin [557]. Key [278]. Kicks [660]. Kidney [519, 165]. Kiefer [675]. Kielcolt [472]. Kinetic [221]. King [619, 810, 749]. Kingdom [722]. Kish [703]. Kitagawa [698]. Klaus [811]. Klockers [755]. Known [377, 210]. Kogan [665]. Kolchin [64]. Kolmogorov [535]. Konig [241]. Kotz [301]. Kpodekpo [216]. Kraeme [753]. Krippendorff [673]. Krishnaiah [437, 479, 475]. Krishnan [16]. Kruse [811]. Kuh [447]. Kurtosis [309, 609].

L [216, 788, 725, 703, 674, 619, 501, 563, 622, 481, 786, 392, 346, 156, 440, 30, 727, 147, 435, 102, 217, 342, 702, 477, 509, 705, 510, 707, 293, 735]. Laboratory [174, 523]. Lamborn [21]. Land [268]. Land-Based [268]. Lane [728]. Language [495]. Lapin [140, 293]. Large [218, 423, 150, 654]. Large-Scale [150]. Larry [510]. Larson [280]. Laser [788]. Laszlo [221]. Latent [735]. Laude [788]. Laura [472]. Law [567, 65]. Lawrence [140, 293]. Leach [20]. Lead [316, 317]. Learned [315]. Least [44, 607, 639]. Left [553]. Legal [169]. Leonard [21, 345]. Leroy [791, 756]. Lesley [770]. Letter [154, 23, 463, 215]. Letters [80, 115, 94, 137]. Level [347]. Levels [315]. Levin [19, 631]. Lewandowski [350]. Lewi [387]. Lewis [102, 441, 582, 99]. Lewis-Beck [582]. Library [219]. Liebetrau [349]. Life [164, 43, 499, 576, 69, 714]. Lifetimes [176]. Light [424, 385]. Likelihood [400, 547, 544, 693, 739, 270, 274, 695, 649, 489, 375]. Limited [780, 526]. Lincoln [436, 296]. Linear [414, 560, 785, 549, 812, 173, 619, 70, 531, 274, 188, 153, 640, 757, 529, 660, 658, 713, 639, 434, 442, 502, 348, 700]. Linearity [417]. Lines [658]. Linhart [708]. Lists [720]. Literature [624]. Litigation [621]. Litter [682, 747]. Litter-Depletion [682, 747]. Little [733]. Load [540, 185, 431, 255]. Load-Curves [540]. Location [778, 330, 543]. Log [544, 693, 739]. Logic [583]. Logit [86, 120, 453, 434]. London [694]. Long [43, 100]. Lonner [731]. Look [179]. Loss [717]. Lothar [624]. Lovie [771]. Low [315, 334]. Lung [366]. Lury [93]. Lynne [383].

M [567, 350, 430, 633, 432, 144, 216, 500, 671, 768, 788, 437, 712, 235, 17, 570, 558, 501, 664, 219, 698, 669, 101, 440, 773, 483, 474, 476, 756, 752, 711, 278, 701, 629, 63, 82, 300, 737, 714, 770, 65, 574, 96, 349, 8, 340, 51]. MacNeill [734]. Macro [258]. Macroeconomic [343]. Mahon [504]. Mail [381]. Majchrzak [430]. Major [350].

Majority [777]. Making [106]. Makridakis [350]. Malaria [277]. Malcolm [749]. Mallows [707]. Malone [725]. Management [322, 244, 599, 770, 722, 754, 701, 631]. Manocha [440]. Manpower [7, 427]. Mansfield [498]. Manual [493]. Manufacturing [373, 799]. Many [753, 528]. Marcia [497]. Marginal [780, 377, 170]. Marginality [13]. Margo [813]. Mark [627, 346, 754, 770]. Marker [172]. Market [284, 428]. Markets [803, 26]. Martin [238]. Mask [50, 119]. Massachusetts [167]. Mastering [482]. Matched [494, 269, 413]. Matérn [620]. Mathai [63, 65]. Mathematical [239, 155, 767, 19, 763, 764, 146, 98]. Mathematicians [228]. Mathematics [634, 449, 783, 618, 347, 766]. Matrices [812]. Matter [9, 24, 28, 41, 4, 25, 32, 35, 47, 62, 78, 104, 116, 52, 57, 76, 87, 110, 127, 148, 159, 201, 118, 129, 160, 202, 211, 249, 265, 213, 232, 266, 276, 290, 310, 273, 302, 307, 325, 339, 360, 371, 313, 326, 351, 361, 396, 420, 450, 464, 485, 372, 397, 421, 451, 465, 512, 541, 585, 604, 487, 542, 586, 587, 637, 652, 678, 716, 605, 638, 679, 680, 740, 774, 792, 816, 741, 742, 793, 794]. Matthew [730]. Maurice [247, 484]. Maximisation [402]. Maximum [547, 544, 693, 739, 270, 274]. Maxwell [619, 562]. Mayes [220]. McCullagh [787]. McCutcheon [735]. McGrath [446]. MDS [219]. Mean [15, 210]. Meaning [140]. Means [21, 773]. Measure [88, 128, 535, 776, 775, 471]. Measured [657]. Measurement [207, 594, 558, 38, 481, 16, 617]. Measurements [643]. Measures [214, 349]. Measuring [780, 309]. Medhi [282, 299]. Medical [275, 762, 285, 59, 796]. Medicine [207]. Megalithic [22]. Membership [790]. Mental [335, 379, 748]. Merits [547]. Merkhofer [701]. Meta [430, 500].

Meta-Analysis [500]. Meta-Analytic [430]. Metabolism [763]. Method [207, 806, 610, 212, 84, 455, 168, 205, 796]. Methodology [731, 186, 71]. Methods [184, 430, 196, 659, 547, 81, 448, 18, 484, 731, 621, 536, 787, 77, 139, 802, 557, 737, 445, 624, 513, 348, 430, 140, 350, 666, 479, 582, 628, 126]. Metric [547]. Meyer [811]. Mezzich [126]. Michael [31, 506, 582]. Micro [258]. Microcomputer [283, 467, 622, 482, 448]. Microcomputers [368]. Middle [6]. Mignot [501]. Mike [584]. Mikel [442, 621]. Miles [68]. Miller [478]. Minimally [173]. Minimising [422]. Minimum [229]. MINITAB [435]. Minor [303]. Misconceived [422]. Missed [268]. Missing [733, 121, 534, 569]. Mitchell [224, 226]. Mitrinović [773]. Mixture [142, 270, 690]. Model [414, 648, 549, 548, 173, 274, 86, 182, 153, 168, 253, 682, 747, 199, 407, 660, 616, 399, 797, 405, 181, 661, 619, 708]. Model-Building [616]. Model-Formulation [548]. Model-Free [747]. Modelling [149, 400, 419, 113, 424, 333, 401, 412, 561, 646, 402, 263, 427, 415, 403, 418, 379, 399, 413, 581]. Models [408, 520, 404, 699, 422, 155, 551, 780, 227, 647, 31, 537, 519, 673, 85, 531, 398, 156, 188, 723, 690, 114, 197, 185, 529, 431, 434, 409, 521, 411, 170, 406, 528, 410, 348, 700, 560, 785, 617, 447, 343]. Moderation [610]. Modern [392, 293, 508]. Molecular [768, 545]. Moments [553, 308, 514]. Monetary [798]. Monitoring [212, 264, 165, 654, 285]. Monstrous [228]. Monte [718, 533, 61, 186, 406, 528, 190, 566]. Montgomery [384]. Moore [709]. Morbidity [379]. Morgan [671]. Morsels [98]. Mortality [694]. Moses [296]. Mosteller [438, 295, 296]. Motion [506]. Motivating [289]. Mould [762]. Moving [187]. Much [315]. Muhammad [294]. Muirhead [279]. Multi [639]. Multi-Step [639]. Multicollinearity [443]. Multidimensional [219, 278, 346, 629]. Multilevel [699]. Multinomial [522, 651, 195, 717, 380]. Multiple [683, 795, 755, 696, 477]. Multiplicative [411]. Multivariate [589, 635, 353, 549, 687, 124, 303, 279, 600, 538, 814, 627, 391, 562, 705, 387]. Munier [712]. Murphy [577]. Murray [125, 579]. Muthu [246]. Myra [504].

N

[633, 706, 807, 16, 752, 277, 728, 738, 616, 445]. N. [30]. Naiman [556]. Namboodiri [16]. Nathan [472]. National [49, 48, 379, 322, 323, 1, 92, 333]. Naturalistic [436]. Naylor [344]. NCB [253]. Need [409]. Neese [447]. Negative [555, 358]. Negotiation [712]. Neil [730]. Nelson [434]. Nested [779]. Networks [566]. Neurotransmitter [364]. Newbold [632]. News [772]. Newton [350]. Nice [449]. Nicholas [390]. Nigam [789]. Nigerian [287]. No [448, 510]. Noise [775]. Noisy [644]. Nomograms [493]. Non [414, 417, 718, 173, 20, 334, 531, 535, 799, 802, 724, 205, 61, 685, 722, 639, 95]. Non-Linear [414, 173, 531, 639]. Non-Linearity [417]. Non-Manufacturing [799]. Non-Normal [718]. Non-Normality [205]. Non-Parametric [20, 535, 724, 685, 95]. Non-Random [61]. Non-Response [802, 722]. Non-Specific [334]. Nonnegative [812]. Nonparametric [479]. Nonparametrics [222]. Nonparametrics-Invariance [222]. Nonresponse [696]. Normal [493, 718, 270, 132, 386, 688, 355, 65, 135, 210]. Normality [205, 684]. Norman [156, 139]. Norpoth [676]. North [671, 256]. Norval [345]. Nosed [119]. Note [655, 11, 719, 496, 657, 89, 453, 267]. Notice [338]. Nuclear [257, 260]. Number [195, 272, 131]. Numbers [61, 807]. Numerical [138, 536, 225, 758, 521, 513]. Numérique [501]. Nutrition [365].

O [560, 665]. Oakes [81]. Obituary [93, 311, 247]. Objective [377]. Observation [644, 453]. Observation-Specific [453]. Observational [296]. Observations [121, 391]. Observed [208]. Obtained [491]. Obtaining [693, 739, 555]. Offering [6]. Office [198]. Officers [14]. Offshore [263]. Oil [264, 258]. Oliveira [243]. One [454]. OPCS [39]. Open [601]. Operational [709, 333]. Operations [263]. Opinion [180, 6]. Opportunity [654]. Optimal [181, 796]. Optimality [289]. Optimisation [547, 474, 237]. Optimization [566, 236]. Option [6]. Ord [710]. Order [329, 583]. Ordering [353]. Organisation [654]. Organometallic [768]. Orientations [204]. Origin [44]. Ornithologists [509]. Orthogonal [607]. Osborne [474]. Ostle [725]. O’Sullivan [766]. Oswald [754]. Other [63, 14, 194]. Our [373]. Outlier [756]. Outliers [441, 99]. Ovulation [192]. Owen [508]. Own [344].

P [216, 475, 769, 671, 241, 437, 709, 732, 219, 439, 505, 773, 476, 752, 278, 787, 479, 789, 625, 728, 64, 616, 782, 564, 222, 571, 83, 672]. Package [467, 444]. Packages [44]. Page [291]. Pain [334]. Paired [806, 123]. Paired-Samples [123]. Pairs [494, 269]. Panel [38]. Panels [593]. Paolo [343]. Paper [472, 484, 434, 433]. Parabolic [50]. Paradoxes [764]. Parallel [803]. Parameters [778, 270, 170]. Parametric [20, 535, 724, 685, 95]. Parasite [178]. Part [19]. Partial [489]. Parts [683]. Parzen [350]. Passage [410]. Passenger [405]. Patel [386]. Patients [165]. Patricia [790]. Patrick [471]. Pattern [19, 73]. Paul [242, 632, 387]. Paula [510]. PC [782].

Pearce [297]. Pearson [630]. Pearsonian [553]. Pecile [763]. Pederzoli [65]. Percentage [469]. Perform [14]. Performance [198, 775, 126]. Period [117]. Perspective [250]. Perspectives [516]. Pertinent [624]. Peter [144, 575, 447, 503, 756, 751, 790, 236, 237]. Pharmaceutical [525]. Pharmacokinetic [221]. Pharmacokinetics [763]. Pharmacological [149]. Phase [256]. Phi [492]. Philip [448]. Philippou [807]. Phillip [628]. Photosensitivity [108]. Physical [274]. Piecewise [470]. Pillemer [385]. Plain [504]. Plane [785]. Planning [408, 250, 321, 642, 7, 296]. Plate [674]. Platek [752]. Plews [17]. Plot [288]. Plots [130]. Plotting [355]. Plus [640]. PM [183]. Pneumoconiosis [332]. Pocock [359]. Pohl [345]. Point [192, 241, 425]. Points [599, 233]. Policy [430, 254]. Politics [484]. Polkinghorne [511]. Pollard [138]. Pollock [143]. Pooling [426]. Pools [208]. Poorly [657]. Population [353, 37, 19, 724, 369, 786, 805]. Populations [778, 318, 1, 92, 315, 112, 49, 48]. Positions [355]. Possible [260, 406, 528, 59]. Possibly [644]. Post [608]. Post-Code [608]. Posterior [191, 168, 406, 528, 514, 190]. Posters [403]. Power [753, 234, 288, 260, 75]. Powerful [186]. Practical [338, 547, 641, 518, 53, 30, 801, 359, 513, 654]. Practice [423, 109, 466, 161, 554, 645, 297, 387]. Practices [748]. Practitioner [720]. Prakasa [727, 217]. Pre [745]. Pre-Sampling [745]. Precisions [455]. Predict [332]. Predicted [658]. Prediction [192, 644, 114, 685, 639, 661]. Predictions [535]. Predictive [380]. Predictor [131]. Predictors [640]. Preliminary [338, 613, 136]. Prentice [342, 147]. Prescreening [194]. Prescribed [275].

Presence [188]. Present [662]. Presentation [169, 97]. Price [54, 272, 803]. Prices [2]. Primary [101]. Primer [341, 623]. Principal [365, 303, 289, 12, 580]. Principle [490]. Principles [222]. Prior [655, 183, 173, 416]. Priors [531]. Probabilities [801, 717, 377, 294]. Probability [778, 767, 125, 615, 95, 355, 153, 555, 82, 434, 764, 280, 65, 130, 308, 729, 67, 471, 750, 564, 246]. Probit [434]. Problem [192, 330, 179, 151, 534, 612, 810, 795, 611, 761]. Problems [373, 517, 671, 187, 516, 618, 333, 125, 507, 531, 304, 199, 178, 738, 722, 186, 704]. Problems-Sequential [507]. Procedure [647, 231, 600, 555, 206]. Procedures [745, 430, 809, 188]. Proceedings [169]. Process [329, 446]. Processes [559, 241, 238, 217, 300, 282, 299]. Processing [66, 765]. Procurement [589, 635, 814]. Product [663]. Production [304]. Professional [29]. Prognostic [214]. Program [721]. Programme [37, 177]. Programming [146, 502, 236, 237, 713]. Programs [219]. Progress [317, 513]. Project [601, 287]. Projecting [640]. Projection [151]. Proof [689]. Propagation [527]. Properties [545]. Proportional [800]. Proportions [378, 492, 452, 75, 134]. Proposal [459]. Prospects [38]. Province [256]. Proxy [607, 150]. Pseudo [175]. Pseudo-Bayesian [175]. Psychological [320]. Psychology [771]. Public [254]. Published [382]. Pugachev [706]. Punch [484]. Puri [789]. Purposes [646].

Quadratic [717]. Qualitative [673, 484]. Quality [384, 597, 590, 592, 150, 600, 601, 685, 571, 595, 602, 328]. Quantifying [642]. Quantile [609]. Quantitative [448, 393, 434, 14, 433, 510]. Quantitiative [472]. Quantity [495]. Quantum [511]. Questioning [489]. Questionnaire [572].

Questions [785, 572, 6]. Queueing

[483, 566]. Queues [615, 241]. Quirk [683]. Quotes [245, 245].

R [350, 239, 21, 499, 704, 475, 769, 666, 437, 712, 478, 733, 125, 669, 392, 480, 809, 630, 559, 279, 474, 476, 750, 95, 752, 711, 147, 479, 670, 697, 477, 63, 502, 762, 385, 393, 705, 433, 616, 294, 676, 713, 96, 804, 700, 224, 226]. R. [499, 69]. Radiation [336, 318, 315]. Raj [791]. Rand [809]. Random [493, 61, 564, 798, 133, 64]. Randomised [324]. Randomization [91]. Range [100]. Rank [209]. Ranking [227]. Ranks [439]. Rao [769, 752, 437, 570, 727, 217]. Rapid [653]. Rapidly [423]. Rate [337, 428]. Rates [538]. Ratio [801, 695, 649, 135, 375]. Rationality [634]. Ratios [120]. Raymond [242]. Read [386]. Readings [705]. Ready [800]. Real [570, 568]. Reasoning [760]. Recapture [659]. Receptors [364]. Recoverable [256]. Red [402]. Reducing [36]. Reduction [341]. Rees [473]. Reference [655, 219, 531, 272]. Regiment [228]. Regina [156]. Regional [157]. Regions [687]. Regression [327, 683, 286, 470, 462, 656, 44, 744, 72, 756, 374, 114, 233, 800, 529, 491, 477, 658, 434, 521, 510, 410, 70, 757, 141]. Regressions [657, 53]. Reign [810]. Reitsch [726]. Rejoinder [30]. Related [461, 348]. Relation [180, 109]. Relationship [274]. Relationships [136, 412, 416]. Relative [691, 546, 636]. Releasing [337]. Reliability [191, 196, 674, 456, 343, 593, 661, 576, 759, 571]. Remains [22]. Remedies [382]. Rendall [235]. Renewable [259]. Report [596]. Representations [120]. Reprise [92]. Republic [14]. Require [14]. Requirements [259]. Rescuing [373]. Research [430, 174, 725, 240, 446, 630, 484, 731, 151, 582, 255, 524, 51, 430, 699, 703, 709, 753, 790, 770, 445, 348]. Reserves [256]. Residential [177]. Residual [329].

Residuals [549]. Resigno [763]. Resources [674]. Response [6, 381, 802, 616, 722, 36]. Responses [783]. Result [289]. Results [454, 150, 379, 383]. Retail [2, 54, 26]. Reuven [566]. Review [140, 142, 567, 350, 710, 430, 5, 504, 633, 784, 438, 772, 560, 627, 699, 785, 239, 21, 384, 432, 761, 144, 66, 146, 155, 499, 704, 216, 500, 471, 388, 441, 577, 475, 578, 730, 767, 769, 806, 812, 783, 297, 766, 671, 241, 768, 788, 666, 442, 617, 781, 220, 341, 725, 576, 482, 703, 437, 572, 674, 712, 706, 235, 242, 511, 17, 570, 575, 31, 218, 447, 558, 619, 709, 281, 298, 573, 223, 291, 292, 501, 568, 618, 345, 478, 563, 708, 733, 753, 732, 664, 668, 219, 807, 20, 81, 97]. Review [98, 124, 125, 141, 240, 729, 698, 16, 584, 622, 755, 244, 579, 481, 70, 734, 786, 19, 138, 448, 507, 669, 673, 143, 443, 446, 392, 480, 111, 68, 101, 726, 749, 809, 389, 18, 346, 677, 630, 559, 69, 156, 238, 386, 439, 440, 505, 561, 565, 615, 620, 773, 623, 483, 632, 279, 503, 506, 383, 474, 581, 472, 484, 731, 754, 805, 813, 476, 580, 473, 665, 675, 756, 750, 95, 566, 696, 727, 752, 771, 711, 278, 147, 763, 621, 810, 765, 634, 344, 435, 102, 787, 390, 7, 67, 139, 217, 221, 243]. Review [391, 479, 562, 342, 670, 556, 702, 757, 701, 343, 751, 557, 629, 316, 22, 157, 569, 697, 789, 477, 626, 431, 790, 225, 277, 63, 82, 300, 737, 498, 502, 625, 714, 762, 770, 385, 393, 434, 509, 764, 728, 280, 508, 582, 736, 705, 759, 758, 282, 299, 808, 359, 64, 65, 738, 444, 436, 449, 574, 301, 382, 433, 616, 760, 782, 99, 294, 296, 445, 624, 667, 676, 236, 237, 631, 811, 510, 564, 628, 713, 222, 245, 791, 96, 295, 571, 583, 707, 394, 804, 293, 349, 497, 8, 340, 83, 126, 246]. Review [145, 51, 672, 100, 387, 348, 347, 735, 700, 224, 226, 763]. Revised [136]. Revision [574]. Rhythms [107]. Richard [432, 223, 383, 95, 631]. Ripley [578]. Risk [720, 318, 674, 342, 701]. Risks [410]. Road [184, 732]. Robbe [672]. Robert [430, 389, 556, 629]. Robust [458, 330, 686, 295, 476, 756]. Robustness [234]. Roebuck [574]. Rohlf [670]. Roland [240]. Role [314, 29]. Roles [790]. Ron [508]. Ronald [785, 141, 383, 758]. Ronchetti [476]. Rosenblatt [82]. Rosenfeld [556]. Rosenthal [430]. Ross [98, 342, 300]. Rossi [575]. Rotation [27]. Rough [800]. Rough-and-Ready [800]. Rounding [529, 308]. Rousseuw [476, 756]. Routine [212, 663]. Row [779]. Row-Column [779]. Roy [311]. Rubin [733, 696]. Rubinstein [566]. Rudin [568]. Rudolf [811]. Rules [180]. Runs [60]. Ryan [435, 435].

S [142, 567, 350, 146, 577, 769, 297, 341, 576, 706, 292, 753, 97, 19, 143, 773, 727, 102, 217, 562, 342, 757, 22, 697, 582, 444, 436, 791, 246, 700]. Sachs [624]. Safety [260]. Sakamoto [698]. Sales [648, 416]. Sales-Advertising [416]. Sample [522, 692, 651, 378, 330, 644, 552, 375, 75, 796]. Samples [718, 123, 663]. Sampling [522, 651, 745, 589, 635, 814, 662, 608, 720, 493, 37, 16, 122, 367, 552, 801, 203, 614, 722, 369, 596, 797, 514, 91, 176, 798, 394, 294]. Samprit [757]. Samuel [301]. Sanford [70]. Sarndal [752]. Sax [755]. Saxena [63]. Scale [543, 150]. Scaling [88, 219, 629, 346, 278]. Scan [291]. Scheme [167]. Schemes [799, 119]. Schmid [97, 97]. Schmidt [241]. Schoepfle [754, 754]. School [610]. School-Based [610]. Schrodt [448]. Schroeder [510]. Schuëller [737]. Science [577, 516, 534, 701, 126, 59]. Science-Comparative [126]. Sciences [448, 607, 472, 434, 433, 510, 20, 809, 771, 714]. Scientist [499, 69]. Scientists [448, 582]. Scott [100]. Sea [256]. Search [704]. Searle [700]. Seasonal [640, 73]. Seber [391]. Second [141, 379]. Secondary [430, 472]. Section [426]. Sedransk [769]. Selecting [778, 786, 600]. Selection [520, 404, 353, 59, 708]. Self [647, 543].

Self-Starting [543]. Selmes [677]. Semi [50]. Semi-Parabolic [50]. Sen [479, 222]. Sensitivity [447, 566, 418, 757]. Sequential [650, 58, 507, 233, 176, 18, 557, 222]. Serial [172]. Serially [656]. Series [350, 430, 475, 437, 417, 448, 484, 731, 536, 646, 426, 260, 643, 73, 444, 581, 667]. Service [648, 49, 48, 322, 323, 1, 92, 333]. Services [86, 335]. Sets [406, 218]. Sevast’yanov [64]. Several [686, 452]. Severity [331]. Shakun [712]. Shaped [381]. Shapes [438]. Shapiro [205]. Sharon [81]. Sheldon [300]. Sherlock [574]. Shocks [415]. Shop [597]. Shop-Floor [597]. Shorack [559]. SI [125]. Sib [555]. Siegel [781]. Sieve [798]. Sign [494, 269, 743]. Significance [466]. Silverman [565]. Similarities [73]. Simon [634]. Simple [212, 591, 658, 724, 205, 797, 514, 661]. Simulation [327, 566, 457, 61, 578]. Simulations [627]. Simultaneous [685]. Sine [378, 75]. Singh [752]. Single [644]. Sinha [576]. Sinitsyn [706]. Sinozuka [737]. Sir [311, 247]. Sire [491]. Six [56]. Size [470, 552, 495, 724, 75, 796]. Sizes [522, 692, 651, 378]. Sjoquist [510]. Skew [190]. Skewness [719, 309]. Skills [14, 677]. Skin [108, 107]. Skorohod [808]. Slater [810]. Slow [187]. Small [718, 469, 321, 663, 752]. Smallest [778]. Smith [225, 760, 445]. Smoking [366]. Snee [383]. Snub [119]. Snub-Nosed [119]. Social [430, 699, 223, 20, 448, 68, 809, 607, 472, 813, 771, 390, 701, 714, 434, 582, 433, 445, 510, 51, 39, 348, 216]. Software [539, 592, 182, 368, 661, 267, 622]. Sokal [670]. Sol [672]. Solomon [126]. Solution [724]. Solutions [671, 664]. Solving [761]. Some [365, 589, 635, 814, 373, 15, 641, 286, 330, 333, 519, 610, 53, 398, 304, 150, 214, 799, 802, 199, 178, 123, 166, 49, 521, 461, 48, 379, 399, 186, 707, 243]. Soren [615]. Sources [430, 235, 805]. South [645]. Span [363].

Spatial [408, 620]. Special [772, 219, 272]. Specific [334, 453]. Specification [153, 619]. Specifications [526]. Specified [134]. Specifying [416]. Spectator [432]. Spectral [667]. Speculative [284]. Spheres [292]. Spiegel [125]. Spline [540]. Split [288]. Split-Plot [288]. Sport [432]. Spreadsheet [746]. Spreadsheets [357]. Sprent [751]. SQC [588]. Square [229, 606, 468]. Squared [452, 75]. Squares [15, 44, 607, 639]. Srivastava [440]. Stage [423]. Stahel [476]. Stainton [102]. Standard [85]. Standardized [572]. Standards [595]. Starting [543, 690]. Startup [223]. Stations [260]. Statistic [546, 636, 468]. Statistical [149, 400, 373, 627, 525, 384, 441, 261, 297, 746, 193, 442, 337, 169, 281, 298, 573, 753, 610, 3, 42, 305, 622, 138, 56, 809, 29, 279, 597, 367, 675, 727, 147, 763, 646, 810, 402, 102, 178, 105, 554, 625, 14, 280, 758, 433, 99, 624, 628, 106, 222, 429, 340, 405, 354, 368, 348, 328, 267, 560, 703, 235, 733, 81, 481, 238, 439, 621, 7, 139, 217, 760]. Statistician [761, 275]. Statisticians [591]. Statistics [710, 338, 314, 174, 283, 216, 598, 388, 475, 766, 454, 341, 482, 437, 242, 558, 330, 345, 20, 125, 109, 244, 10, 212, 121, 250, 46, 257, 74, 200, 565, 805, 476, 161, 590, 752, 771, 479, 702, 469, 800, 63, 498, 714, 457, 166, 782, 513, 357, 321, 287, 224, 226, 21, 146, 730, 767, 769, 17, 223, 478, 732, 141, 698, 579, 392, 68, 18, 559, 503, 473, 750, 787, 243, 556, 557, 762, 764, 574, 8, 567, 432, 671, 666, 781, 725, 292, 240, 632, 770, 631, 811, 804, 293, 347]. Statistics. [140]. Statististics [509]. Step [639]. Stephan [510]. Stephen [558, 581]. Stereology [239]. Stewart [430, 618]. Steyn [697]. Stigler [558]. Stochastic [746, 570, 505, 506, 217, 236, 237, 578, 706, 734, 389, 561, 665, 300, 737, 282, 299, 808]. Stock [261, 262]. Stoodley [102, 388]. Stopping [180]. Storage [665]. Stoyanov [729]. Strata [204]. Strategic [253].

Strategies [208]. Strategy [284]. Stream [261]. Stronger [289]. Structural [673, 737, 447]. Structure [404]. Structures [356, 410]. Stuart [710, 359, 394]. Student [577]. Studies [184, 207, 318, 537, 19, 723, 316, 413, 766, 81, 244, 296, 19, 82]. Study [414, 430, 113, 152, 718, 107, 456, 657, 677, 460, 150, 317, 258, 722]. Stumpf [697]. Subjective [180]. Subjects [753]. Substrates [687]. Success [599]. Suggested [208]. Sum [683, 799]. Summing [385]. SuperCalc [664]. Support [322, 323, 672]. Suppression [374]. Surface [768, 768]. Surfaces [616]. Survey [720, 467, 38, 456, 262, 72, 472, 460, 151, 765, 722, 39, 379, 572, 16, 480, 630]. Surveys [55, 268, 122, 722, 36, 593, 71, 696, 390]. Survival [530, 537, 114, 206, 156]. Suspect [534]. Swedish [262]. Symmetry [719]. Symposium [752]. System [191, 648, 322, 196, 264, 367, 344]. Systematic [430, 662, 575, 526, 754, 754]. Systems [323, 404, 777, 305, 527, 672, 706, 506, 157, 736, 808, 672, 505]. Szekely [764].

T [146, 766, 671, 223, 732, 580, 102, 139, 277, 738, 449, 83, 672]. Tables [438, 550, 120, 230, 459, 377, 786, 225]. Tail [132, 26]. Takkenberg [672]. Tales [56]. Tapia [95]. Tashman [21]. Tatham [705]. Taxonomy [126]. Taylor [581, 497]. Teaching [283, 275, 10, 46, 503, 597, 89, 457, 357]. Tech [601]. Technique [423]. Techniques [725, 31, 456, 7, 252, 235, 138, 749, 102, 508]. Technologies [259]. Technology [669]. Temperature [107]. Tendencies [208]. Tensor [787]. Terms [306]. Test [164, 454, 494, 330, 269, 459, 695, 649, 123, 205, 743]. Testing [681, 415, 663, 26, 576]. Tests [613, 466, 60, 30, 532, 684, 461, 375, 285, 354, 363]. Textbook [341]. Textbooks [123]. Texts [278]. Thalassaemia [194]. Thanks [11].

Their [331, 630, 773, 14]. Them [88].

Theorem [777]. Theory [710, 785, 155, 297, 673, 466, 161, 727, 634, 629, 157, 764, 280, 256, 710, 483, 279, 82, 281, 298, 125]. There [121, 528]. Thieman [753]. Thisted [758]. Thom [22, 22]. Thomas [141, 439, 435]. Thompson [95]. Thoughts [399]. Three [553, 77]. Threshold [647, 695, 649]. Tiago [243]. Tijms [561]. Time [350, 475, 437, 417, 581, 147, 536, 646, 426, 366, 643, 73, 667, 236, 237, 410, 444]. Times [198]. Timing [285]. Toby [441, 99]. Toit [697]. Tolerance [687]. Tool [746]. Tools [582]. Topics [812]. Total [90]. Tracking [428]. Trade [772, 306]. Trading [258, 595]. Traffic [405]. Training [592, 14, 722]. Transformation [530, 378]. Transformations [496]. Transplant [165]. Transplantation [519]. Treating [177]. Treatise [440]. Treatment [686, 534]. Treatments [108]. Tree [189]. Trends [438, 694, 640]. Trial [117, 105]. Trials [525, 108, 180, 109, 111, 30, 359]. Trouble [123]. Trout [383]. Truncation [490]. TSPACK [444]. Tufte [393]. Tukey [438, 295]. Tumour [172]. Tumour-Marker [172]. Two [117, 378, 270, 330, 550, 456, 423, 539, 60, 120, 658, 135, 75, 134, 181]. Two-Period [117]. Two-Sample [330]. Two-Stage [423]. Two-Way [550, 120, 181]. Type [121, 613].

U [241]. U.K [379]. U.K. [748]. UK

[805, 261, 136]. Ultimate [256]. UMP [743]. Umphrey [734]. Unbalanced [700]. Uncertainty [558, 642]. Undergraduate [287]. Understanding [630, 556, 751, 510, 295]. Underwriting [197]. Unemployment [694, 319, 537, 320, 206]. Unequal [552, 801, 294]. Uniformity [684]. Unifying [490]. Unimodal [462]. United [722]. Units [798]. Univariate [647, 270]. University [287]. Unknown [425].

Unsound [533]. Unusual [233]. Urban [157]. Usage [331, 264]. Use [608, 180, 744, 229, 121, 253, 469, 489, 488, 61, 411, 416]. Used [182, 214]. Useful [398]. User [219, 5]. Using [336, 283, 548, 378, 107, 303, 657, 72, 553, 367, 332, 402, 377, 538, 434, 186, 528, 514, 267, 382].

V [784, 239, 767, 241, 706, 217, 789, 225, 625, 808, 64, 119]. V-Mask [119]. Vague [811]. Validation [334]. Validity [151, 446]. Value [530, 778]. Values [690, 658]. Vandaele [81]. Vapnik [301]. Variability [642]. Variable [547, 780, 657]. Variables [551, 607, 60, 600, 135, 461, 564]. Variance [662, 15, 234, 214, 491, 488, 170, 12, 562, 676]. Variances [523]. Variates [493]. Variation [210, 620]. Vasić [773]. Versus [133]. Vic [441, 99]. Victimization [38]. Victims [177]. Viertl [750]. View [319]. Visual [283, 393]. Vladimir [64, 301]. Vol [430, 754]. Volodin [738]. Volume [710, 40, 103, 158, 248, 312, 370, 486, 603, 715, 815, 731, 459, 236, 237]. Volumes [402]. Vote [584]. Vries [672]. Vulnerability [320].

W [430, 438, 144, 447, 708, 81, 392, 630, 565, 731, 476, 702, 701, 629, 697, 431, 449, 294, 96, 295]. W. [769]. Walsh [713]. Walter [568, 81, 731]. Wards [694]. Water [424, 674]. Watson [292, 121]. Wautelet [788]. Way [550, 120, 181]. Wayne [617, 622]. Ways [21]. Weather [263]. Weekes [623]. Wegener [704]. Weighting [352]. Weights [88]. Weisberg [81, 70, 145]. Weiss [730]. Wellner [559]. Werner [754]. Wetherill [557, 18]. Whales [268]. Where [121]. Whittaker [223]. Whittemore [342]. Whittle [505, 236, 237]. Who [14]. Whole [683]. Wider [206]. Wilcox [809]. William [582, 296]. Williams [83]. Wilson [157]. Winkler [350]. Winters [641]. Within [92, 1, 49, 48]. Without [657]. Wolf [500, 235]. Women [192]. Wonnacott [141, 141]. Woodland [90]. Woodward [622, 770]. Work [287, 39]. Workers [725]. World [511]. Wrong [398].

X [219].

Yadalah [146]. Yadolah [569]. Yin [430]. Young [629]. Youngman [51]. Yvonna [436].

Z [759]. Zellner [655]. Zero [85]. Zirkel [556]. Zucchini [708]. Zupan [218]. Zurbenko [667].

References

Gandy:1979:CCP

[1] R. J. Gandy. The calculation of catchment populations within the National Health Service. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):29–37, March 1979. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                             www.jstor.org/stable/2988030. See

comments [48].

Carruthers:1980:RDR

[2] A. G. Carruthers, D. J. Sellwood, and P. W. Ward. Recent developments in the retail prices index. Journal of the Royal Statistical Society. Series D (The Statistician), 29(1):1–32, March 1980. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                             www.jstor.org/stable/2987492.

Greenfield:1980:SCB

[3] Tony Greenfield and Sean Siday. Statistical computing for business and industry. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (1):33–55, March 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987493.

Anonymous:1980:FMa

[4] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(1):56, March 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987491.

Armstrong:1980:URG

[5] William Armstrong. A User’s review of GLIM3. Journal of the Royal Statistical Society. Series D (The Statistician), 29(1):57–63, March 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987494.

Kalton:1980:EOM

[6] G. Kalton, Julie Roberts, and D. Holt. The effects of offering a middle response option with opinion questions. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (1):65–78, March 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987495.

Morgan:1980:BRB

[7] Byron J. T. Morgan. Book review: Statistical Techniques for Manpower Planning, by D. J. Bartholomew; A. F. Forbes. Journal of the Royal Statistical Society. Series D (The Statistician), 29(1):79, March 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987496.

Thomas:1980:BRB

[8] D. A. H. Thomas. Book review: A Basic Course in Statistics, by G. M. Clarke; D. Cooke. Journal of the Royal Statistical Society. Series D (The Statistician), 29(1):79–80, March 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987497.

Anonymous:1980:BMa

[9] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(1):??, March 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987498.

Hawkes:1980:TEA

[10] A. G. Hawkes. Teaching and examining applied statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):81–89, June 1980. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988109.

Anonymous:1980:NT

[11] Anonymous. A note of thanks. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):90, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988110.

Smith:1980:GCP

[12] R. H. Smith. A geometric comparison of principal component analysis with analysis of variance. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):91–96, June 1980. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988111.

Preece:1980:CAF

[13] D. A. Preece. Covariance analysis, factorial experiments and marginality. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (2):97–122, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988112.

Richards:1980:TSO

[14] Winston A. Richards. Training of statistical officers and other officers of the Republic who require quantitative skills to perform their duties effectively. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2): 123–134, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988113.

Boyd:1980:EMS

[15] K. T. Boyd. The expectation of mean squares in analysis of variance — some conjectures. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):135–142, June 1980. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988114.

Harris:1980:BRB

[16] P. T. Harris. Book review: Survey Sampling and Measurement, by N. Krishnan Namboodiri. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):143, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988115.

Ellison:1980:BRB

[17] R. Ellison. Book review: Introductory Statistics, by A. M. Plews. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):143–144, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988116.

Jeffers:1980:BRB

[18] J. N. R. Jeffers. Book review: Sequential Methods in Statistics, by G. B. Wetherill. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):144, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988117.

Hawkes:1980:BRB

[19] Alan G. Hawkes. Book review: Studies in Mathematical Biology — Part I: Cellular Behaviour and the Development of Pattern; Studies in Mathematical Biology — Part II: Population and Communities, by S. A. Levin. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):144–145, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988118.

Greenfield:1980:BRB

[20] Tony Greenfield. Book review: Introduction to Statistics: A Non-Parametric Approach for The Social Sciences, by Chris Leach. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):145, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988119.

Bissell:1980:BRB

[21] Derek Bissell. Book review: The Ways and Means of Statistics, by Leonard J. Tashman; Kathleen R. Lamborn. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (2):146, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988120.

Porteous:1980:BRB

[22] Hugh L. Porteous. Book review: Megalithic Remains in Britain and Brittany, by A. Thom; A. S. Thom. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):146–149, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988121.

Haybittle:1980:LE

[23] J. L. Haybittle and L. S. Freedman. Letter to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):149, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988122.

Anonymous:1980:BMb

[24] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):??, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988123.

Anonymous:1980:FMb

[25] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(2):??, June 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988108.

Williams:1980:TRM

[26] A. G. Prys Williams and R. W. E. Wadman. Testing retail markets: A cautionary tail. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (3):151–165, September 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988039.

Darton:1980:RFA

[27] R. A. Darton. Rotation in factor analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 29(3): 167–194, September 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988040.

Anonymous:1980:BMc

[28] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(3):195– 204, September 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988044.

Jones:1980:RPE

[29] J. D. Jones and G. K. Kanji. The role of professional experience in statistical education. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (3):196–203, September 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988041.

Liddell:1980:PTC

[30] Douglas Liddell. Practical tests for comparative trials: A rejoinder to N. L. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (3):205–207, September 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988042.

Ethell:1980:BRB

[31] Bernard Ethell. Book review: Econometric Models, Techniques, and Applications, by Michael D. Intriligator. Journal of the Royal Statistical Society. Series D (The Statistician), 29(3): 208, September 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988043.

Anonymous:1980:FMc

[32] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(3):??, September 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988038.

Kish:1980:DED

[33] Leslie Kish. Design and estimation for domains. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (4):209–222, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987728.

Bailar:1980:EDC

[34] Barbara A. Bailar and Charles D. Jones. The evaluation of the 1980 decennial census. Journal of the Royal Statistical Society. Series D (The Statistician), 29 (4):223–235, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987729.

Anonymous:1980:FMd

[35] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(4):236, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987727.

Sudman:1980:RRE

[36] Seymour Sudman. Reducing response error in surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 29(4):237–273, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987730.

Fellegi:1980:EPC

[37] Ivan P. Fellegi. Evaluation programme of the 1976 census of population and housing — a sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 29(4):275–312, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987731.

Fienberg:1980:MCV

[38] S. E. Fienberg. The measurement of crime victimization: Prospects for panel analysis of a panel survey. Journal of the Royal Statistical Society. Series D (The Statistician), 29(4):313– 350, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987732.

Whitehead:1980:WSS

[39] F. E. Whitehead. The work of the social survey division of OPCS. Journal of the Royal Statistical Society. Series D (The Statistician), 29(4):351– 363, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987733.

Anonymous:1980:VI

[40] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 29(4):364, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987726.

Anonymous:1980:BMd

[41] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 29(4):??, December 1980. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987734.

Hand:1981:BBS

[42] D. J. Hand. Branch and bound in statistical data analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):1–13, March 1981. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987699.

Bissinger:1981:LLA

[43] Barnard H. Bissinger. A long life analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (1):15–21, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987700.

Gordon:1981:ECP

[44] H. A. Gordon. Errors in computer packages. least squares regression through the origin. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):23–29, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987701.

Preece:1981:DFD

[45] D. A. Preece. Distributions of final digits in data. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):31–60, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987702.

Idil:1981:DTS

[46] Orhan Idil. The dilemma of teaching statistics and applications to business decisions in the developing countries. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (1):61–67, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987703.

Anonymous:1981:BMa

[47] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):68–76, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987708.

Skidmore:1981:SCC

[48] Steve Skidmore. Some comments on “The Calculation of Catchment Populations within the National Health Service” by Gandy. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):69–72, March 1981. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987704. See

[1].

Senn:1981:SCC

[49] Stephen J. Senn. Some comments on “The Calculation of Catchment Populations within the National Health Service”. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (1):73–75, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987705.

Bissell:1981:CSP

[50] A. F. Bissell. Correction to “A Semi-Parabolic Mask for Cusum Charts”. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (1):77, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987706.

Wallace:1981:BRB

[51] W. Wallace. Book review: Analysing Social and Educational Research Data, by M. B. Youngman. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):78, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987707.

Anonymous:1981:FMa

[52] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):??, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987698.

Jeffers:1981:IAR

[53] J. N. R. Jeffers. Investigation of alternative regressions: Some practical examples. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):79–88, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987560.

Morgan:1981:ERP

[54] D. D. V. Morgan. Estimation of retail price changes. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):89–96, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987561.

Fessey:1981:BCS

[55] M. C. Fessey. Business censuses and surveys in developing countries. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):97–106, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987562.

Hunter:1981:SST

[56] William G. Hunter. Six statistical tales. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2): 107–117, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987563.

Anonymous:1981:FMb

[57] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):118–130, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987559.

Gadsden:1981:SAA

[58] R. J. Gadsden and G. K. Kanji. Sequential analysis for angular data. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):119–129, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987564.

Wood:1981:PSE

[59] P. D. P. Wood. A possible selection effect in medical science. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):131–135, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987565.

Jolliffe:1981:RTD

[60] I. T. Jolliffe. Runs tests for detecting dependence between two variables. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2): 137–141, June 1981. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987566.

Sculli:1981:UNR

[61] D. Sculli. The use of non-random numbers in Monte Carlo simulation. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2): 143–151, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987567.

Anonymous:1981:BMb

[62] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):152, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987576.

Rao:1981:BRBa

[63] M. Bhaskara Rao. Book review: The H-Function with Applications in Statistics and Other Disciplines, by A. M. Mathai; R. K. Saxena. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):153, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987568.

Shanbhag:1981:BRBa

[64] D. N. Shanbhag. Book review: Random Allocations, by V. F. Kolchin; B. A. Sevast’yanov; Vladimir P. Chistyakov. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (2):153, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987569.

Shanbhag:1981:BRBb

[65] D. N. Shanbhag. Book review: Characterizations of the Normal Probability Law, by A. M. Mathai; G. Pederzoli. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (2):153–154, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987570.

Brown:1981:BRB

[66] P. S. Brown. Book review: Introduction to Business Data Processing, by D. K. Carver. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):154–155, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987571.

Morgan:1981:BRB

[67] Byron J. T. Morgan. Book review: An Introduction to Applied Probability, by I. F. Blake. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):155–156, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987572.

Imber:1981:BRBa

[68] Valerie Imber. Book review: Demystifying Social Statistics, by John Irvine; Ian Miles; Jeff Evans. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):156–157, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987573.

Kanji:1981:BRB

[69] G. K. Kanji. Book review: R. A. Fisher, The Life of a Scientist, by Joan Fisher Box. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (2):157–158, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987574.

Hassard:1981:BRB

[70] T. H. Hassard. Book review: Applied Linear Regression, by Sanford Weisberg. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):158, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987575.

Wilson:1981:IMD

[71] Paul Wilson. Improving the methodology of drinking surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):159– 167, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988046.

Holt:1981:RAU

[72] D. Holt and A. J. Scott. Regression analysis using survey data. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):169– 178, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988047.

Raveh:1981:GDC

[73] Adi Raveh. A graphic display for characterization of seasonal pattern similarities of time series. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):179–192, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988048.

Kanji:1981:DAS

[74] G. K. Kanji and W. G. Gilchrist. A degree in applied statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):193– 198, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988049.

Ury:1981:CCA

[75] Hans K. Ury. Continuity-corrected approximations to sample size or power when comparing two proportions: Chisquared or arc sine? Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):199–203, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988050.

Anonymous:1981:FMc

[76] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):204, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988045.

Morgan:1981:TAM

[77] Byron J. T. Morgan. Three applications of methods of cluster-analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):205– 223, September 1981. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988051.

Anonymous:1981:BMc

[78] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):224– 230, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988057.

Poloniecki:1981:FAH

[79] J. Poloniecki and D. Daniel. Further analysis of the hills and armitage enuresis data. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (3):225–229, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988052.

Beale:1981:LE

[80] E. M. L. Beale, P. W. Lane, J. A. Nelder, A. F. Bissell, and P. L. Goldsmith. Letters to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):231–234, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988053.

Greenfield:1981:BRBa

[81] Tony Greenfield. Book review: Statistical Methods for Comparative Studies, by Sharon Anderson; Ariane Auquier; Walter W. Hauck; David Oakes; Walter Vandaele; Herbert I. Weisberg. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):235, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988054.

Rao:1981:BRBb

[82] M. Bhaskara Rao. Book review: Studies in Probability Theory, by M. Rosenblatt. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (3):236, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988055.

Thunhurst:1981:BRB

[83] Colin Thunhurst. Book review: Drink in Great Britain 1900–1979, by G. P. Williams; G. T. Brake. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):236– 238, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988056.

Jinkinson:1981:CDG

[84] R. A. Jinkinson and M. Slater. Critical discussion of a graphical method for identifying discrete distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):239– 248, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988060.

Hietikko:1981:EZI

[85] Harri Hietikko. On the effect of zero initial estimates on the standard estimation of ARMA (p, q) models. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):249– 254, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988061.

Kanji:1981:LMA

[86] G. K. Kanji. Logit model to assess the environmental impacts caused by bus services. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):255–261, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988062.

Anonymous:1981:FMd

[87] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):262– 296, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988059.

Adams:1981:MWE

[88] Robert M. Adams and John Wilmut. A measure of the weights of examination components, and scaling to adjust them. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (4):263–269, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988063.

Morrison:1981:NTA

[89] S. J. Morrison. A note on the teaching of AoV. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (4):271–274, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988064.

Rennolls:1981:TAW

[90] Keith Rennolls. The total area of woodland in Berkshire is …. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):275– 287, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988065.

Walters:1981:SRD

[91] D. E. Walters. Sampling the randomization distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):289–295, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988066.

Gandy:1981:CCP

[92] R. J. Gandy. The calculation of catchment populations within the National Health Service — Reprise. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):297– 302, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988067. See [1, 48].

Anonymous:1981:ODA

[93] Anonymous. Obituary: Dennis Albert Lury, 1922–81. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):303–304, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988068.

Haws:1981:LE

[94] A. P. Haws and H. A. Gordon. Letters to Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (4):304–308, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988069.

Leonard:1981:BRB

[95] Tom Leonard. Book review: Non-Parametric Probability Density Estimation, by Richard A. Tapia; James R. Thompson. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):309–310, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988070.

Stone:1981:BRB

[96] Mervyn Stone. Book review: Discrete Discriminant Analysis, by M. Goldstein; W. R. Dillon. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):311, December 1981. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988071.

Greenfield:1981:BRBb

[97] Tony Greenfield. Book review: Handbook of Graphic Presentation, by C. F. Schmid; S. E. Schmid. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):311– 312, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988072.

Greenfield:1981:BRBc

[98] Tony Greenfield. Book review: Mathematical Morsels, by Ross Honsberger. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):

312–313, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988073.

Siday:1981:BRB

[99] S. E. Siday. Book review: Outliers in Statistical Data, by Vic Barnett; Toby Lewis. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (4):313–315, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988074.

Ward:1981:BRB

[100] R. Ward. Book review: Long Range Forecasting (From Crystal Ball to Computer), by J. Scott Armstrong. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):315, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988075.

Imber:1981:BRBb

[101] Valerie Imber. Book review: Evaluating Primary Care, by Ewen M. Clark; J. A. Forbes. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (4):316, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988076.

Merrett:1981:BRB

[102] J. D. Merrett. Book review: Applied Statistical Techniques, by K. D. C. Stoodley; T. Lewis; C. L. S. Stainton. Journal of the Royal Statistical Society. Series D (The Statistician), 30 (4):316, December 1981. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988077.

Anonymous:1981:VI

[103] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):317– 318, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988058.

Anonymous:1981:BMd

[104] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):??, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988078.

Pocock:1981:SAC

[105] Stuart J. Pocock. Statistical aspects of clinical trial design. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):1–18, March 1981. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988097.

Spiegelhalter:1981:SAC

[106] D. J. Spiegelhalter. Statistical aids in clinical decision-making. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):19–36, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988098.

Dunstan:1981:FSE

[107] F. D. J. Dunstan, S. Barham, K. W. Kemp, A. B. J. Nix, R. J. Rowlands, D. W. Wilson, M. J. Phillips, and K. Griffiths. A feasibility study for early detection of breast cancer using breast skin temperature rhythms. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (1):37–52, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988099.

Corbett:1981:CCT

[108] M. F. Corbett, A. Davis, J. L. M. Hawk, A. Herxheimer, I. A. Magnus, M. Mohindra, N. J. Tan, and P. Toosi. Controlled clinical trials of treatments for skin photosensitivity. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):53–62, March 1981. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988100.

Grimshaw:1981:SRC

[109] J. J. Grimshaw and G. V. Jaffe. Statistics in relation to clinical trials in general practice. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):63–69, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988101.

Anonymous:1981:FMe

[110] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):70, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988096.

Huitson:1981:RCT

[111] A. Huitson, J. Poloniecki, R. Hews, and N. Barker. A review of cross-over trials. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (1):71–80, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988102.

Senn:1981:EHC

[112] S. J. Senn and W. B. Samson. Estimating hospital catchment populations. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (1):81–96, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988103.

Campbell:1981:CSD

[113] M. G. Campbell and S. M. Rudolfer. A comparative study of diagnostic modelling in electrodiagnosis. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):97–106, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988104.

OQuigley:1981:RMS

[114] John O’Quigley. Regression models and survival prediction. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):107–116, March 1981. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988105.

Cressie:1981:LE

[115] Noel Cressie. Letters to Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):117, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988106.

Anonymous:1981:BMe

[116] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(1):118, March 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988107.

Armitage:1982:TPC

[117] P. Armitage and M. Hills. The twoperiod crossover trial. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2):119–131, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987883.

Anonymous:1982:FMa

[118] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2):132–168, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987882.

Rowlands:1982:SNV

[119] R. J. Rowlands, A. B. J. Nix, M. A.

Abdollahian, and Kenneth W. Kemp. Snub-nosed V-mask control schemes. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2): 133–142, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987884.

Karger:1982:ALA

[120] J. C. Karger. An application of logit analysis to the effective graphical representations of two-way tables of ratios. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (2):143–148, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987885.

Honohan:1982:UDW

[121] Patrick Honohan and Colm McCarthy. On the use of Durbin–Watson type statistics where there are missing observations. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (2):149–152, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987886.

Kiregyera:1982:SFA

[122] Ben Kiregyera. On sampling frames in African censuses and surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2):153–167, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987887.

Preece:1982:TTC

[123] D. A. Preece. t is for trouble (and textbooks): A critique of some examples of the paired-samples t-test. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2):169–195, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987888.

Greenfield:1982:BRBa

[124] Tony Greenfield. Book review: Introduction to Multivariate Analysis, by Christopher Chatfield; Alexander J. Collins. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2):196, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987889.

Greenfield:1982:BRBb

[125] Tony Greenfield. Book review: Theory and Problems of Probability and Statistics, SI, by Murray R. Spiegel. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2):196, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987890.

Thunhurst:1982:BRB

[126] Colin Thunhurst. Book review:

Taxonomy and Behavioral Science-Comparative Performance of Grouping Methods, by J. E. Mezzich; H. Solomon. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2): 197–198, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987891.

Anonymous:1982:BMa

[127] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(2):??, June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987892.

Armstrong:1982:CDM

[128] William Armstrong. A chart to describe and measure the effects of industrial inspection errors. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):199–209, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987986.

Anonymous:1982:FMb

[129] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):210– 230, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987985.

Stirling:1982:EAI

[130] W. Douglas Stirling. Enhancements to aid interpretation of probability plots. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):211– 220, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987987.

Yeomans:1982:GKC

[131] Keith A. Yeomans and Paul A. Golder. The Guttman–Kaiser criterion as a predictor of the number of common factors. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3): 221–229, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987988.

Hawkes:1982:ANT

[132] Alan G. Hawkes. Approximating the normal tail. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):231–236, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987989.

Wynn:1982:CVR

[133] H. P. Wynn. Controlled versus random experimentation. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):237–244, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987990.

Ury:1982:CTP

[134] Hans K. Ury. Comparing two proportions: Finding p_{2} when p_{1}, N, α, and β are specified. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):245–250, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987991.

Shanmugalingam:1982:ART

[135] S. Shanmugalingam. On the analysis of the ratio of two correlated normal variables. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (3):251–258, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987992.

Holden:1982:RBP

[136] K. Holden and D. A. Peel. The relationships between preliminary and revised estimates of GDP and consumers’ expenditure in the UK. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):259– 266, September 1982. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987993.

Mullins:1982:LE

[137] Eamonn Mullins and David Griffiths. Letters to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):267–270, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987994.

Hawkes:1982:BRB

[138] Alan G. Hawkes. Book review: A Handbook of Numerical and Statistical Techniques, by J. H. Pollard. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):271, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987995.

Morgan:1982:BRB

[139] G. W. Morgan. Book review: Statistical Methods in Biology, by Norman T. J. Bailey. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (3):272, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987996.

Agha:1982:BRBa

[140] M. Agha. Book review: Statistics. Meaning and Methods, by Lawrence Lapin. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (3):272–273, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987997.

Greenfield:1982:BRBc

[141] Tony Greenfield. Book review: Regression: A Second Course in Statistics, by Thomas H. Wonnacott; Ronald J. Wonnacott. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (3):273–274, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987998.

Agha:1982:BRBb

[142] M. Agha. Book review: Finite Mixture Distribution, by B. S. Everitt; D. J. Hand. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (3):274, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987999.

Hewitt:1982:BRB

[143] V. N. Hewitt. Book review: The Algebra of Econometrics, by D. S. G. Pollock. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3): 274–275, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988000.

Boyd:1982:BRB

[144] K. T. Boyd. Book review: Incomplete Block Designs, by Peter W. M. John. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3): 275, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988001.

Vaughan:1982:BRB

[145] T. B. Vaughan. Book review: An Introduction to Data Analysis, by B. D. Bowen; H. F. Weisberg. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):275– 277, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988002.

Burley:1982:BRBa

[146] H. T. Burley. Book review: Mathematical Programming in Statistics, by T. S. Arthanari; Yadalah Dodge. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):277, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988003.

Mackenzie:1982:BRB

[147] Gilbert Mackenzie. Book review: The Statistical Analysis of Failure Time Data, by J. D. Kalbfleisch; R. L. Prentice. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (3):278, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988004.

Anonymous:1982:BMb

[148] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):??, September 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988005.

Adena:1982:SMP

[149] Michael A. Adena. Statistical modelling of a pharmacological experiment. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4): 279–291, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987630.

Martin:1982:QPI

[150] Jean Martin and Bob Butcher. The quality of proxy information — some results from a large-scale study. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4):293– 319, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987631.

Lightman:1982:IPV

[151] Ernie S. Lightman. Indirect projection and the validity problem in survey research. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (4):321–331, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987632.

Clarke:1982:CSE

[152] G. M. Clarke and S. A. Ricketts. A case study in the efficiency of experimental design. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (4):333–337, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987633.

Missiakoulis:1982:ASL

[153] Spyros Missiakoulis. An alternative specification of the linear probability model. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (4):339–343, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987634.

Cohen:1982:LE

[154] Ayala Cohen. Letter to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4): 345–346, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987635.

Burley:1982:BRBb

[155] H. T. Burley. Book review: Game Theory, Mathematical Models of Conflict, by A. J. Jones. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4):347, December 1982. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987636.

Kay:1982:BRB

[156] Richard Kay. Book review: Survival Models and Data Analysis, by Regina C. Elandt-Johnson; Norman L. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 31 (4):347, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987637.

Porteous:1982:BRB

[157] H. L. Porteous. Book review: Catastrophe Theory and Bifurcation-Applications to Urban and Regional Systems, by A. G. Wilson. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4):348, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987638.

Anonymous:1982:VI

[158] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4):349– 350, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987628.

Anonymous:1982:BMc

[159] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4):??, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987639.

Anonymous:1982:FMc

[160] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 31(4):??, December 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987629.

Lindley:1982:TPB

[161] Dennis V. Lindley. Theory and practice of Bayesian statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):1– 11, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987587.

DeGroot:1982:CEF

[162] Morris H. DeGroot and Stephen E. Fienberg. The comparison and evaluation of forecasters. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):12–22, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987588.

Zellner:1982:ABA

[163] Arnold Zellner. Applications of Bayesian analysis in econometrics. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):23– 34, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987589.

Barlow:1982:EIL

[164] Richard E. Barlow and Jaw Huan Hsiung. Expected information from a life test experiment. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):35–45, March/ June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987590.

Smith:1982:MKT

[165] A. F. M. Smith, M. West, K. Gordon, M. S. Knapp, and I. M. G. Trimble. Monitoring kidney transplant patients. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):46– 54, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987591.

Rubin:1982:SAB

[166] Donald B. Rubin. Some applications of Bayesian statistics to educational data. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):55– 68, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987592.

Dumouchel:1982:MAI

[167] William H. Dumouchel. The 1982 Massachusetts Automobile Insurance Classification Scheme. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):69–81, March/ June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987593.

Naylor:1982:CMC

[168] J. C. Naylor and A. F. M. Smith. A contamination model in clinical chemistry: An illustration of a method for the efficient computation of posterior distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (1/2):82–87, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987594.

Fienberg:1982:PBS

[169] Stephen E. Fienberg and Joseph B.

Kadane. The presentation of Bayesian statistical analyses in legal proceedings. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):88– 98, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987595.

Skene:1982:CMD

[170] A. M. Skene. Computing marginal distributions for the dispersion parameters of analysis of variance models. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):99– 108, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987596.

Smith:1982:FAC

[171] J. Q. Smith. Forecasting accident claims for an assurance company. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):109– 115, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987597.

Tate:1982:BAA

[172] Helen Tate. A Bayesian approach to the analysis of serial tumour-marker data. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2): 116, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987598.

Eaves:1982:MIP

[173] David M. Eaves. Minimally informative prior analysis of a non-linear model. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2): 117, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987599.

Baker:1982:BSI

[174] A. G. Baker and D. R. Lane-Joynt.

Bayesian statistics in an industrial research laboratory. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):118–123, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987600.

Andrews:1982:PBB

[175] R. W. Andrews and T. M. F. Smith.

Pseudo-Bayesian and Bayesian approach to auditing. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):124–126, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987601.

Wright:1982:SBA

[176] David E. Wright. Sequential batch acceptance sampling for exponential lifetimes (extended abstract). Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):127– 130, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987602.

Schervish:1982:RPT

[177] Mark J. Schervish and Elizabeth Elmer.

A residential programme for treating victims of child abuse. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):131–137, March/ June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987603.

Piccinato:1982:SSP

[178] Ludovico Piccinato. Some statistical problems in parasite experiments. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):138– 143, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987604.

Kokolakis:1982:NLP

[179] G. E. Kokolakis. A new look at the problem of classification with binary data. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2): 144–152, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987605.

Freedman:1982:ASO

[180] L. S. Freedman and D. J. Spiegelhalter. The assessment of the subjective opinion and its use in relation to stopping rules for clinical trials. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):153– 160, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987606.

Verdinelli:1982:CBO

[181] Isabella Verdinelli. Computing Bayes D- and A-optimal block designs for a two-way model. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):161–167, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987607.

Meinhold:1982:BAC

[182] Richard J. Meinhold and Nozer D.

Singpurwalla. Bayesian analysis of a commonly used model for describing software failures. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):168–173, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987608.

Chaloner:1982:ABP

[183] Kathryn M. Chaloner and George T.

Duncan. Assessment of a beta prior distribution: PM elicitation. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):174– 180, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987609.

Abbess:1982:BMA

[184] C. Abbess, D. Jarrett, and C. C.

Wright. Bayesian methods applied to road accident blackspot studies. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):181– 187, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987610.

Polasek:1982:HLC

[185] Wolfgang Polasek. Hierarchical load curve models. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):188–194, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987611.

Stewart:1982:BAU

[186] Leland Stewart. Bayesian analysis using Monte Carlo integration — a powerful methodology for handling some difficult problems. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):195–200, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987612.

DeWit:1982:IPS

[187] J. R. De Wit. Inventory problems with slow moving items: A Bayesian approach. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (1/2):201–206, March/June 1982. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987613.

Makov:1982:ABP

[188] Udi E. Makov. Approximate Bayesian procedures for dynamic linear models in the presence of jumps. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):207– 213, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987614.

Houle:1982:GTB

[189] Alfred Houle. The genealogical tree of Bayesians. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):214–215, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987615.

vanDijk:1982:MCA

[190] H. K. van Dijk and T. Kloek. Monte Carlo analysis of skew posterior distributions: An illustrative econometric example. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (1/2):216–223, March/June 1982. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987616.

Abdel-Wahed:1982:APD

[191] A. R. Abdel-Wahed and Alan Winterbottom. Approximating posterior distributions of system reliability. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):224– 228, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987617.

Carter:1982:BCP

[192] Randy L. Carter and Barry J. N. Blight.

A Bayesian change point problem with an application to the prediction and detection of ovulation in women. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):229– 230, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987618.

Conrad:1982:DSE

[193] S. A. Conrad. The design of statistical experiments — a computational approach. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (1/2):231–232, March/June 1982. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987619.

Steve:1982:PBT

[194] G. Steve and A. Rossi-Mori. Prescreening of beta-thalassaemia carriers: A comparison between Bayesian and other approaches. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):233–239, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987620.

Boender:1982:BAN

[195] C. G. E. Boender and A. H. G. Rinnooy Kan. A Bayesian analysis of the number of cells of a multinomial distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (1/2):240–248, March/June 1982. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987621.

Bacon-Shone:1982:BMS

[196] John Bacon-Shone. Bayesian methods for system reliability and availability. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2): 249–251, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987622.

Philips:1982:BMC

[197] Lawrence D. Philips and Thomas K. Wisniewski. Bayesian models for computer-aided underwriting. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):252– 263, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987623.

Burridge:1982:JTO

[198] J. Burridge and A. O’Hagan. Job times and Office performance. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):264– 272, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987624.

Pericchi:1982:SPB

[199] L. R. Pericchi and I. Rodriguez-Iturbe.

On some problems in Bayesian model choice in hydrology. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):273–278, March/ June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987625.

Kanji:1982:BAS

[200] Gopal K. Kanji. Bulletin in Applied Statistics (BIASs). Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):279, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987626.

Anonymous:1982:BMd

[201] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2): ??, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987627.

Anonymous:1982:FMd

[202] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 32(1/2):

??, March/June 1982. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987586.

Moors:1983:BES

[203] J. J. A. Moors. Bayes’ estimation in sampling for auditing. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):281–288, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987933.

Gadsden:1983:ACS

[204] R. J. Gadsden and G. K. Kanji. Analysis of clay strata orientations. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):289– 296, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987934.

Royston:1983:SME

[205] J. P. Royston. A simple method for evaluating the Shapiro–Francia W′ test of non-normality. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):297–300, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987935.

Royston:1983:WAS

[206] G. H. D. Royston. Wider application of survival analysis: An evaluation of an unemployment benefit procedure. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):301– 306, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987936.

Altman:1983:MMA

[207] D. G. Altman and J. M. Bland. Measurement in medicine: The analysis of method comparison studies. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):307– 317, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987937.

Stefani:1983:OBT

[208] Raymond T. Stefani. Observed betting tendencies and suggested betting strategies for European football pools. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):319– 329, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987938.

Lawson:1983:RAC

[209] Andrew Lawson. Rank analysis of covariance: Alternative approaches. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):331– 337, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987939.

Sinha:1983:BEM

[210] S. K. Sinha. Bayesian estimation of the mean of a normal distribution when the coefficient of variation is known. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):339– 345, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987940.

Anonymous:1983:BMa

[211] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):346, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987955.

Healy:1983:SMM

[212] M. J. R. Healy. A simple method for monitoring routine statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):347– 349, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987941.

Anonymous:1983:FMa

[213] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):350, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987932.

OQuigley:1983:AVF

[214] John O’Quigley. Asymptotic variance formulae for some measures used as prognostic indices. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):351–352, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987942.

Townsend:1983:LE

[215] W. Townsend. Letter to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3): 353, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987943.

Casimir:1983:BRB

[216] J. F. R. Casimir. Book review: Social and Economic Statistics for Africa, by G. M. C. Kpodekpo; P. L. Arya. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3): 354, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987944.

Morgan:1983:BRBa

[217] G. W. Morgan. Book review: Statistical Inference for Stochastic Processes, by I. V. Basawa; B. L. S. Prakasa Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):354– 355, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987945.

Everitt:1983:BRB

[218] B. S. Everitt. Book review: Clustering of Large Data Sets, by Jure Zupan. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (3):355, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987946.

Gordon:1983:BRB

[219] A. D. Gordon. Book review: The User’s Guide to Multidimensional Scaling, with Special Reference to the MDS(X) Library of Computer Programs, by A.

P. M. Coxon; P. M. Davies. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):355– 356, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987947.

Cormack:1983:BRB

[220] R. M. Cormack. Book review: Applications of Econometrics, by D. G. Mayes. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (3):356, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987948.

Morgan:1983:BRBb

[221] G. W. Morgan. Book review: Kinetic Data Analysis, Design and Analysis of Enzyme and Pharmacokinetic Experiments, by Laszlo Endrenyi. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):356– 357, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987949.

Sprent:1983:BRBa

[222] P. Sprent. Book review: Sequential Nonparametrics-Invariance Principles and Statistical Inference, by P. K. Sen. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (3):357, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987950.

Fuller:1983:BRB

[223] Michael F. Fuller. Book review: Introducing Social Statistics, by Richard Startup; Elwyn T. Whittaker. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):357– 358, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987951.

leRoux:1983:BRBa

[224] A. A. le Roux. Book review: European Historical Statistics, 1750–1970, by B. R. Mitchell. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):358–359, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987952.

Raggett:1983:BRB

[225] G. F. Raggett. Book review: Tables for Numerical Integration, by H. V. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3): 359, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987953.

leRoux:1983:BRBb

[226] A. A. le Roux. Book review: International Historical Statistics: Africa and Asia, by B. R. Mitchell. Journal of the Royal Statistical Society. Series D (The Statistician), 32(3):359– 360, September 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987954.

Cohen:1983:AGF

[227] Ayala Cohen and C. L. Mallows. Assessing goodness of fit of ranking models to data. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (4):361–374, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987538.

Pearce:1983:MRM

[228] S. C. Pearce. The monstrous regiment of mathematicians. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):375–378, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987539.

Harris:1983:UMC

[229] R. R. Harris and G. K. Kanji. On the use of minimum chi-square estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4): 379–394, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987540.

Kenett:1983:EAC

[230] Ron S. Kenett. On an exploratory analysis of contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):395– 403, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987541.

Holmes:1983:GIP

[231] D. I. Holmes. A graphical identification procedure for growth curves. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):405– 415, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987542.

Anonymous:1983:FMb

[232] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):416, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987537.

Paul:1983:SDU

[233] S. R. Paul. Sequential detection of unusual points in regression. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):417– 424, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987543.

Kanji:1983:FIR

[234] G. K. Kanji and C. K. Liu. Further investigation of robustness of power in the analysis of variance. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):425– 438, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987544.

Edmond:1983:BRB

[235] W. J. Edmond. Book review: Statistical Sources and Techniques, by F. J. Rendall; D. M. Wolf. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):439, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987545.

Smith:1983:BRBa

[236] David K. Smith. Book review: Optimization Over Time. Volume 1. Dynamic Programming and Stochastic Control, by Peter Whittle. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):439, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987546.

Smith:1983:BRBb

[237] David K. Smith. Book review: Optimisation Over Time. Volume 2. Dynamic Programming and Stochastic Control, by Peter Whittle. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):439–440, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987547.

Kay:1983:BRB

[238] Richard Kay. Book review: Statistical Analysis of Counting Processes, by Martin Jacobsen. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):440, December 1983. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987548.

Bingham:1983:BRB

[239] M. S. Bingham. Book review: Combinatorial Integral Geometry with Applications to Mathematical Stereology, by R. V. Ambartzumian. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):440–441, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987549.

Greenfield:1983:BRB

[240] Tony Greenfield. Book review: Statistics in Research and Development, by Roland Caulcutt. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):441–442, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987550.

Coleman:1983:BRB

[241] Rodney Coleman. Book review: Queues and Point Processes, by P. Franken; D. Konig; U. Arndt; V. Schmidt. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4): 442, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987551.

Edwards:1983:BRB

[242] Bernard Edwards. Book review: Basic Statistics for Business and Economics, by Paul G. Hoel; Raymond J. Jessen. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4): 442–443, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987552.

Morgan:1983:BRBc

[243] G. W. Morgan. Book review: Some Recent Advances in Statistics, by J. Tiago de Oliveira; B. Epstein. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):443, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987553.

Hart:1983:BRB

[244] Anna Hart. Book review: Applied Statistics for Management Studies, by David Croft. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (4):443–444, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987554.

Sprent:1983:BRBb

[245] P. Sprent. Book review: Quotes, Damned Quotes, and …, by John Bibby. Journal of the Royal Statistical Society. Series D (The Statistician), 32 (4):444, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987556.

Tricker:1983:BRB

[246] A. R. Tricker. Book review: Probability and Errors, by S. K. Muthu. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4): 444, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987555.

Bartholomew:1983:OSM

[247] D. J. Bartholomew. Obituary: Sir Maurice Kendall FBA. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):445–446, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987557.

Anonymous:1983:VI

[248] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):447– 448, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987536.

Anonymous:1983:BMb

[249] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 32(4):??, December 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987558.

Hutber:1983:ESP

[250] F. W. Hutber. Energy statistics and the planning perspective. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):1–7, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987710.

Lyness:1983:GDF

[251] F. K. Lyness. Gas demand forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (1):9–21, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987711.

Rhys:1983:TFE

[252] J. M. W. Rhys. Techniques for forecasting electricity demand. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):23–33, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987712.

Ormerod:1983:DUN

[253] Richard Ormerod and John McLeod.

The development and use of the NCB strategic model. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):35–49, March 1983. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987713.

Brown:1983:FPP

[254] Philip J. Brown. Forecasts and public policy. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (1):51–63, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987714.

Skinner:1983:LRA

[255] N. H. Skinner. Load research and its application to electricity demand forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):65–73, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987715.

Rosing:1983:EUE

[256] K. E. Rosing and P. R. Odell. Estimating the ultimate economically recoverable hydrocarbon reserves of the North sea province: Phase. 1 the theory. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (1):75–90, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987716.

Jones:1983:SNE

[257] P. M. S. Jones. Statistics and nuclear energy. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (1):91–102, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987717.

Scanlan:1983:MMC

[258] Tony Scanlan. Micro or macro? A case study of approaches to forecasting based on oil trading (1970–74). Journal of the Royal Statistical Society. Series D (The Statistician), 33(1): 103–110, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987718.

Buckley-Golder:1983:CRE

[259] D. H. Buckley-Golder, R. G. Derwent, K. F. Langley, J. F. Walker, and A. V. Ward. Contribution of renewable energy technologies to future energy requirements. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (1):111–132, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987719.

Openshaw:1983:ESC

[260] S. Openshaw. An evaluation of the safety characteristics of current and possible future nuclear power stations series. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (1):133–142, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987720.

Chesshire:1983:SAU

[261] John Chesshire. A statistical analysis of the UK stream boiler stock. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1): 143–150, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987721.

Holgersson:1983:SEC

[262] Margareta Holgersson. A survey of energy consumption in the Swedish housing stock. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):151–160, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987722.

OCarroll:1983:WMO

[263] F. M. O’Carroll. Weather modelling for offshore operations. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):161–169, March 1983. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987723.

Jessup:1983:ACC

[264] Gina Jessup. Avon County Council fuel oil usage monitoring system. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1): 171–179, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987724.

Anonymous:1983:BMc

[265] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):??, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987725.

Anonymous:1983:FMc

[266] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(1):??, March 1983. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987709.

Yeo:1984:NCU

[267] Gee-Kin Yeo. A note of caution on using statistical software. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):181–184, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987848.

Horwood:1984:CMW

[268] J. W. Horwood. Compensating for missed whales in land-based surveys of bowhead whales. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):185–200, June 1984. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987849.

Hay:1984:AST

[269] Alan Hay and Francis Peck. An alternative to the sign test in a matched pairs design. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (2):201–204, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987850.

Everitt:1984:MLE

[270] B. S. Everitt. Maximum likelihood estimation of the parameters in a mixture of two univariate normal distributions; a comparison of different algorithms. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (2):205–215, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987851.

Brown:1984:FBG

[271] Philip Brown and Clive Payne. Forecasting the 1983 British general election. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (2):217–228, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987852.

Silver:1984:CCB

[272] M. S. Silver. Criteria for choosing between alternative consumer price index number formulae with special reference to chained indices. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):229–237, June 1984. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987853.

Anonymous:1984:FMa

[273] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):238, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987847.

Hui:1984:PIM

[274] Y. V. Hui and Y. K. Tse. A physical interpretation of the maximum likelihood estimation of a linear functional relationship model. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):239–242, June 1984. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987854.

Gore:1984:TED

[275] Sheila M. Gore. Teaching experimental design: Prescribed by a medical statistician. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (2):243–247, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987855.

Anonymous:1984:BMa

[276] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):248, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987862.

Raggett:1984:BRB

[277] G. F. Raggett. Book review: The Biomathematics of Malaria, by N. T. J. Bailey. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (2):249–250, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987856.

MacFie:1984:BRB

[278] H. J. H. MacFie. Book review: Key Texts in Multidimensional Scaling, by P. M. Davies; A. P. M. Coxon. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):250–251, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987857.

Kent:1984:BRB

[279] John T. Kent. Book review: Aspects of Multivariate Statistical Theory, by R.

J. Muirhead. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):251, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987858.

Riley:1984:BRB

[280] I. S. Riley. Book review: Introduction to Probability Theory and Statistical Interference, by Harold J. Larson. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):251, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987859.

Freeman:1984:BRBa

[281] G. H. Freeman. Book review: Theory of Statistical Experiments, by H. Heyer. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):252, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987860.

Selby:1984:BRBa

[282] Michael J. P. Selby. Book review: Stochastic Processes, by J. Medhi. Journal of the Royal Statistical Society. Series D (The Statistician), 33(2):252, June 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987861.

Bland:1984:UMV

[283] John Martin Bland. Using a microcomputer as a visual aid in the teaching of statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):253–259, September 1984.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988229.

Tuckwell:1984:CSS

[284] R. H. Tuckwell. Choice of strategy in a speculative market. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):261–271, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988230.

Tate:1984:TCT

[285] Helen Tate. The timing of “confirmatory” tests in medical monitoring. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):273– 281, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988231.

Currie:1984:SCB

[286] Iain Currie and Athol Korabinski. Some comments on bivariate regression. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):283– 293, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988232.

Uche:1984:UPW

[287] P. I. Uche. Undergraduate project work in statistics at a Nigerian university. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):295– 300, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988233.

Kanji:1984:PAS

[288] G. K. Kanji and C. K. Liu. Power aspects of split-plot designs. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):301– 311, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988234.

OHagan:1984:MPC

[289] A. O’Hagan. Motivating principal components, and a stronger optimality result. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3): 313–315, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988235.

Anonymous:1984:BMb

[290] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):316, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988247.

Fuller:1984:BRB

[291] M. F. Fuller. Book review: Page Scan. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):317– 318, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/view/2988236.

Gadsden:1984:BRB

[292] R. J. Gadsden. Book review: Statistics on Spheres, by G. S. Watson. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):318– 319, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988237.

Targett:1984:BRB

[293] David Targett. Book review: Statistics for Modern Business Decisions, by Lawrence L. Lapin. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):319, September 1984. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988238.

Skinner:1984:BRB

[294] C. J. Skinner. Book review: Sampling with Unequal Probabilities, by K. R. W. Brewer; Muhammad Hanif. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):319– 320, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988239.

Stuart:1984:BRB

[295] Michael Stuart. Book review: Understanding Robust and Exploratory Data Analysis, by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):320– 321, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988240.

Smeeton:1984:BRB

[296] Nigel C. Smeeton. Book review:

Planning and Analysis of Observational Studies, by William G. Cochran; Lincoln E. Moses; Frederick Mosteller. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):321– 322, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988241.

Cobby:1984:BRB

[297] J. M. Cobby. Book review: The Agricultural Field Experiment. A Statistical Examination of Theory and Practice, by S. C. Pearce. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):322, September 1984. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988242.

Freeman:1984:BRBb

[298] G. H. Freeman. Book review: Theory of Statistical Experiments, by H. Heyer. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):322– 323, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988243.

Selby:1984:BRBb

[299] Michael J. P. Selby. Book review: Stochastic Processes, by J. Medhi. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3): 323, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988244.

Rao:1984:BRB

[300] M. Bhaskara Rao. Book review: Stochastic Processes, by Sheldon M. Ross. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (3):323–324, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988245.

Shoesmith:1984:BRB

[301] E. Shoesmith. Book review: Estimation of Dependences Based on Empirical Data, by Vladimir Vapnik; Samuel Kotz. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (3):324, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988246.

Anonymous:1984:FMb

[302] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):??, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988228.

Hawkins:1984:EMD

[303] Douglas M. Hawkins and L. Paul Fatti.

Exploring multivariate data using the minor principal components. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4):325– 338, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987737.

Kmietowicz:1984:SPC

[304] Z. W. Kmietowicz and M. J. Al-Fallogi.

Some problems in the calculation of indices of industrial production in developing countries: The case of Iraq, 1954– 73. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4):

339–349, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987738.

Hand:1984:SES

[305] D. J. Hand. Statistical expert systems: Design. Journal of the Royal Statistical Society. Series D (The Statistician), 33 (4):351–369, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987739.

Greenfield:1984:TTI

[306] C. C. Greenfield. The terms of trade index and terms of trade effect. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4):371– 379, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987740.

Anonymous:1984:FMc

[307] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4):380, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987736.

Tricker:1984:ERM

[308] A. R. Tricker. Effects of rounding on the moments of a probability distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4): 381–390, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987741.

Groeneveld:1984:MSK

[309] Richard A. Groeneveld and Glen Meeden. Measuring skewness and kurtosis. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4): 391–399, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987742.

Anonymous:1984:BMc

[310] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4):400, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987744.

Anonymous:1984:OSR

[311] Anonymous. Obituary: Sir Roy Allen, CBE, FBA (1906–1983). Journal of the Royal Statistical Society. Series D (The Statistician), 33(4):401–406, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987743.

Anonymous:1984:VI

[312] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 33(4):407, December 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987735.

Anonymous:1985:FMa

[313] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):1–2, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987499.

Bailey:1985:RSC

[314] Norman T. J. Bailey. The role of statistics in controlling and eradicating infectious diseases. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):3–17, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987500.

Gilbert:1985:HMC

[315] E. S. Gilbert. How much can be learned from populations exposed to low levels of radiation? Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):19–30, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987501.

Pocock:1985:ELC

[316] Stuart J. Pocock and Deborah Ashby. Environmental lead and Children’s intelligence: A review of recent epidemiological studies. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):31–44, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987502.

Raab:1985:ELS

[317] Gillian M. Raab, Mary Fulton, Duncan P. H. Laxen, and George O. B. Thomson. The Edinburgh lead study: Aspects of design and progress. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):45–57, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987503.

Darby:1985:ERR

[318] Sarah C. Darby. Evaluation of radiation risk from epidemiological studies of populations exposed at high doses. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1): 59–72, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987504.

Cook:1985:CVU

[319] Derek G. Cook. A critical view of the unemployment and health debate. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1): 73–82, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987505.

Jackson:1985:DVP

[320] Paul R. Jackson. Differential vulnerability and psychological health in unemployment. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):83–91, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987506.

Thunhurst:1985:ASA

[321] Colin Thunhurst. The analysis of small area statistics and planning for health. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1): 93–106, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987507.

Ashford:1985:MDS

[322] J. R. Ashford and T. C. Bailey. A management decision support system for the National Health Service. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):107–116, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987508.

Bailey:1985:IDS

[323] T. C. Bailey. Information and decision support systems in the National Health Service. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):117–124, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987509.

Altman:1985:CRG

[324] Douglas G. Altman. Comparability of randomised groups. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):125–136, ???? 1985. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987510.

Anonymous:1985:BMa

[325] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(1):137, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987511.

Anonymous:1985:FMb

[326] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):139–140, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988152.

Beale:1985:RBB

[327] E. M. L. Beale. Regression: A bridge between analysis and simulation. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):141–154, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988153.

Woodall:1985:SDQ

[328] William H. Woodall. The statistical design of quality control charts. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2): 155–160, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988154.

Bora:1985:CDF

[329] A. C. Bora. Change over design with first order residual effects and errors following a first order autoregressive process. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (2):161–173, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988155.

Fung:1985:SRT

[330] Karen Yuen Fung, Hyunshik Lee, and I. Tajuddin. Some robust test statistics for the two-sample location problem. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (2):175–182, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988156.

Hutchinson:1985:ASD

[331] T. P. Hutchinson. Analysing severity data when assessors differ in their usage of the categories. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):183–195, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988157.

Maclaren:1985:UDA

[332] W. M. Maclaren. Using discriminant analysis to predict attacks of complicated pneumoconiosis in coalworkers. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2): 197–208, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988158.

Gilchrist:1985:SAM

[333] Robert Gilchrist. Some aspects of modelling operational problems in the National Health Service. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):209–214, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988159.

Heinrich:1985:VAE

[334] I. Heinrich, H. O’Hare, B. Sweetman, and J. A. D. Anderson. Validation aspects of an empirically derived classification for ‘non-specific’ low back pain. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2): 215–230, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988160.

Marioni:1985:EMH

[335] Alessandra Marioni, Eugenio Torre, and Mario Comelli. Evaluating mental health services in Italy. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):231–235, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988161.

Barry:1985:EIE

[336] S. F. Barry. Estimation of internal exposure to radiation using empirical Bayes. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2): 237–242, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988162.

Faragher:1985:SER

[337] E. B. Faragher, E. N. Chantler, and C. T. Roberts. A statistical evaluation of the rate of degradation of copper releasing IUCDs. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):243–248, ???? 1985. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988163.

Anonymous:1985:PND

[338] Anonymous. Preliminary notice: Dates for your diary. Conference on practical Bayesian statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):249, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988164.

Anonymous:1985:BMb

[339] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):250, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988176.

Thomas:1985:BRB

[340] D. A. H. Thomas. Book review: Basic Statistical Computing, by D. Cooke; A. H. Craven; G. M. Clarke. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):251–252, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988165.

Craggs:1985:BRB

[341] Carolyn Craggs. Book review: A Primer in Data Reduction: An Introductory Statistics Textbook, by A. S. C. Ehrenberg. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (2):252–253, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988166.

Muirhead:1985:BRB

[342] C. R. Muirhead. Book review: Environmental Epidemiology: Risk Assessment, by Ross L. Prentice; Alice S. Whittemore. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (2):253–254, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988167.

Oliver:1985:BRB

[343] F. R. Oliver. Book review: Evaluating the Reliability of Macroeconomic Models, by Gregory C. Chow; Paolo Corsi. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2): 254–255, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988168.

McNair:1985:BRB

[344] Robert McNair. Book review: Build Your Own Expert System, by Chris Naylor. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (2):255, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988169.

Gibson:1985:BRB

[345] M. G. Gibson. Book review: Basic Statistics for Business & Economics, by Leonard J. Kazmier; Norval F. Pohl. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2): 256–257, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988170.

Jeffers:1985:BRB

[346] J. N. R. Jeffers. Book review: Multidimensional Scaling, by Mark L. Davison. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (2):257–258, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988171.

Williamson:1985:BRB

[347] R. J. Williamson. Book review: Statistics for Advanced Level Mathematics, by I. Gwyn Evans. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):258–259, ???? 1985. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988172.

Williams:1985:BRB

[348] A. G. Prys Williams. Book review: Linear Statistical Models and Related Methods; With Applications to Social Research, by John Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):259–260, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988173.

Tarling:1985:BRB

[349] Roger Tarling. Book review: Measures of Association, by Albert M. Liebetrau. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2): 260–261, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988174.

Anderson:1985:BRB

[350] Oliver D. Anderson. Book review: The Forecasting Accuracy of Major Time Series Methods, by S. Makridakis; A. Andersen; R. Carbone; R. Fildes; M. Hibon; R. Lewandowski; J. Newton; E. Parzen; R. Winkler. Journal of the Royal Statistical Society. Series D (The Statistician), 34(2):261–262, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988175.

Anonymous:1985:FMc

[351] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):263–264, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987652.

French:1985:WEC

[352] Simon French. The weighting of examination components. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):265–280, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987653.

Breth:1985:DDP

[353] M. Breth. On data description, population estimation and the selection of multivariate ordering. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):281–285, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987654.

Westberg:1985:CIS

[354] Margareta Westberg. Combining independent statistical tests. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):287–296, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987655.

Looney:1985:PPP

[355] Stephen W. Looney and Thomas R.

Gulledge, Jr. Probability plotting positions and goodness of fit for the normal distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):297–303, ???? 1985. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987656.

Lee:1985:FDC

[356] Sik-Yum Lee and Wai-Yin Poon. Further developments on constrained estimation in analysis of covariance structures. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (3):305–316, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987657.

Soper:1985:STS

[357] Jean B. Soper and Martin P. Lee. Spreadsheets in teaching statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3): 317–321, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987658.

Ross:1985:NBD

[358] G. J. S. Ross and D. A. Preece. The negative binomial distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):323–335, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987659.

Seldrup:1985:BRB

[359] Jorgen Seldrup. Book review: Clinical Trials. A Practical Approach, by Stuart J. Pocock. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):337–338, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987660.

Anonymous:1985:BMc

[360] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(3):??, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987661.

Anonymous:1985:FMd

[361] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):341–342, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987820.

Kanji:1985:E

[362] Gopal K. Kanji. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):343, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987821.

Williamson:1985:STC

[363] R. J. Williamson. Span tests on cusums.

Journal of the Royal Statistical Society. Series D (The Statistician), 34(4): 345–356, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987822.

Dunn:1985:ADN

[364] Graham Dunn. Assays for drug, neurotransmitter and hormone receptors. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4): 357–364, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987823.

Ali:1985:PCA

[365] Asghar Ali, G. M. Clarke, and K. Trustrum. Principal component analysis applied to some data from fruit nutrition experiments. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):365–370, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987824.

Peace:1985:TCB

[366] L. R. Peace. A time correlation between cigarette smoking and lung cancer. Journal of the Royal Statistical Society. Series D (The Statistician), 34 (4):371–381, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987825.

Kriens:1985:SSI

[367] J. Kriens and R. H. Veenstra. Statistical sampling in internal control by using the AOQL-System. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):383–390, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987826.

Wetherill:1985:DES

[368] G. Barrie Wetherill, John B. Curram, J. P. Burman, P. Duncombe, J. E. Hailstone, and R. E. Kempson. The design and evaluation of statistical software for microcomputers. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):391–427, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987827.

Tam:1985:CFP

[369] S. M. Tam. On covariance in finite population sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):429–433, ???? 1985. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987828.

Anonymous:1985:VI

[370] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):435–439, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987819.

Anonymous:1985:BMd

[371] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 34(4):??, ???? 1985. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987829.

Anonymous:1986:FMa

[372] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):1–2, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988292.

Barnard:1986:ROM

[373] G. A. Barnard. Rescuing our manufacturing industry-some of the statistical problems. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):3–16, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988293.

Lewis:1986:SEB

[374] Jerry W. Lewis and Luis A. Escobar. Suppression and enhancement in bivariate regression. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):17–26, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988294.

Shoukri:1986:SLR

[375] M. M. Shoukri. K-sample likelihood ratio tests on the inverse-Gaussian distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):27–32, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988295.

Preece:1986:IEI

[376] D. A. Preece. Illustrative examples: Illustrative of what? Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):33–44, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988296.

Pelz:1986:EPC

[377] Wolfgang Pelz and I. J. Good. Estimating probabilities from contingency tables when the marginal probabilities are known, by using additive objective functions. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (1):45–50, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988297.

Dobson:1986:SSC

[378] Annette J. Dobson and Val J. Gebski.

Sample sizes for comparing two independent proportions using the continuitycorrected arc sine transformation. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):51–53, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988298.

Smeeton:1986:MEM

[379] Nigel C. Smeeton. Modelling episodes of mental illness: Some results from the Second U.K. National Morbidity Survey. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (1):55–63, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988299.

Wood:1986:APC

[380] John B. Wood. An application of predictive checks in the Bayesian analysis of multinomial data. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):65–72, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988300.

McGowan:1986:FSC

[381] Ian McGowan. Fitting S-shaped curves to mail response data. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):73–77, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988301.

Shoesmith:1986:BRBa

[382] E. Shoesmith. Book review: Using Published Data: Errors and Remedies, by Herbert Jacob. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):79–80, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988302.

Kinsella:1986:BRBa

[383] A. Kinsella. Book review: Experiments in Industry: Analysis and Interpretation of Results, by Ronald D. Snee; Lynne B. Hare; J. Richard Trout. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):80–81, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988303.

Bissell:1986:BRBa

[384] Derek Bissell. Book review: Introduction to Statistical Quality Control, by Douglas C. Montgomery. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):81–82, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988304.

Reese:1986:BRBa

[385] Allan Reese. Book review: Summing Up, by R. J. Light; D. B. Pillemer. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1): 82–83, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988305.

Kemp:1986:BRBa

[386] A. W. Kemp. Book review: Handbook of the Normal Distribution, by J. K. Patel; C. B. Read. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):84, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988306.

Wetherill:1986:BRB

[387] G. Barrie Wetherill. Book review: Multivariate Data Analysis in Industrial Practice, by Paul J. Lewi. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):85, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988307.

Chanter:1986:BRBa

[388] D. O. Chanter. Book review: Applied and Computational Statistics; A First Course, by Keith Stoodley. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):85–86, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988308.

Jacka:1986:BRB

[389] S. D. Jacka. Book review: Stochastic Calculus and Applications, by Robert J. Elliott. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):87, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988309.

Moon:1986:BRB

[390] Nick Moon. Book review: Data Construction in Social Surveys, by Nicholas Bateson. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):87–88, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988310.

Morgan:1986:BRBa

[391] D. D. V. Morgan. Book review: Multivariate Observations, by G. A. F. Seber. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (1):88–89, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988311.

Holmes:1986:BRB

[392] P. Holmes. Book review: A Modern Approach to Statistics, by R. L. Iman; W. J. Conover. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):89–90, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988312.

Reese:1986:BRBb

[393] R. Allan Reese. Book review: The Visual Display of Quantitative Information, by Edward R. Tufte. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):90, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988313.

Sugden:1986:BRB

[394] R. A. Sugden. Book review: The Ideas of Sampling, by Alan Stuart. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):91, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988314.

Anonymous:1986:A

[395] Anonymous. Announcement. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):92, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988315.

Anonymous:1986:BMa

[396] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):??, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988316.

Anonymous:1986:FMb

[397] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):93–94, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987512.

Kanji:1986:EAM

[398] G. K. Kanji. Editorial: “All models are wrong but some are useful”. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2): 95–96, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987513.

Smith:1986:SBT

[399] A. F. M. Smith. Some Bayesian thoughts on modelling and model choice.

Journal of the Royal Statistical Society. Series D (The Statistician), 35(2): 97–101, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987514.

Aitkin:1986:SML

[400] Murray Aitkin. Statistical modelling: The likelihood approach. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):103–113, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987515.

Harding:1986:MCA

[401] E. F. Harding. Modelling: The classical approach. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):115–134, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987516.

McLaren:1986:SMD

[402] Christine E. McLaren, Gary M. Brittenham, Victor R. Gordeuk, Margaret A. Hughes, and Louise J. Keating. Statistical modelling of the distribution of red blood cell volumes in iron deficiency anemia using the expectation-maximisation algorithm. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):135–142, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987517.

Sharot:1986:MAP

[403] Trevor Sharot. Modelling the audience to posters. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):143–150, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987518.

Beck:1986:SSM

[404] M. B. Beck. The selection of structure in models of environmental systems. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (2):151–161, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987519.

VanDerLaan:1986:SMC

[405] B. S. Van Der Laan and A. S. Louter. A statistical model for the costs of passenger car traffic accidents. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):163–174, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987520.

Stewart:1986:BPD

[406] Leland Stewart and William W. Davis.

Bayesian posterior distributions over sets of possible models with inferences computed by Monte Carlo integration. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2): 175–182, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987521.

Pettit:1986:DBM

[407] L. I. Pettit. Diagnostics in Bayesian model choice. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):183–190, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987522.

Baxter:1986:GPM

[408] Mike Baxter. Geographical and planning models for data on spatial flows. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2): 191–198, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987523.

Reese:1986:DAN

[409] R. A. Reese. Data analysis: The need for models? Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):199–206, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987524.

Whitmore:1986:FPT

[410] G. A. Whitmore. First-passage-time models for duration data: Regression structures and competing risks. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):207–219, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987525.

Shepherd:1986:UAM

[411] J. G. Shepherd and M. D. Nicholson. Use and abuse of multiplicative models in the analysis of fish catch-at-age data. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (2):221–227, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987526.

Johnston:1986:MEC

[412] F. R. Johnston, P. J. Harrison, A. S.

Marshall, and K. M. France. Modelling and the estimation of changing relationships. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (2):229–235, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987527.

Thompson:1986:MMC

[413] S. G. Thompson. Modelling in matched case-control studies in epidemiology. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2): 237–244, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987528.

Adcock:1986:CLA

[414] C. J. Adcock. Choosing linear approximations to a non-linear model: A case study. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (2):245–249, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987529.

Sander:1986:MTS

[415] Peter C. Sander. Modelling and testing shocks. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (2):251–262, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987530.

Young:1986:UPI

[416] R. J. Young and M. Young. Use of prior information in specifying the dynamics of sales-advertising relationships. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2): 263–269, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987531.

Davies:1986:DNL

[417] Neville Davies and Joseph D. Petruccelli. Detecting non-linearity in time series. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (2):271–280, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987532.

Skene:1986:BMS

[418] A. M. Skene, J. E. H. Shaw, and T. D. Lee. Bayesian modelling and sensitivity analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):281–288, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987533.

Brown:1986:MAA

[419] Gordon Brown. Modelling advertising awareness. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):289–299, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987534.

Anonymous:1986:BMb

[420] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(2):??, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987535.

Anonymous:1986:FMc

[421] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):301–302, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987745.

Bendell:1986:MMM

[422] Tony Bendell. Minimising misconceived models. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):303–309, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987746.

Gilks:1986:PTS

[423] W. R. Gilks, S. E. Hills, and S. M.

Gore. In practice: A two-stage technique for rapidly exploring large datasets. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3): 311–324, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987747.

Carr:1986:MAB

[424] Martin R. Carr. Modelling the attenuation of broad band light down a water column. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):325–333, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987748.

Freeman:1986:UCP

[425] James M. Freeman. An unknown change point and goodness of fit. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):335–344, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987749.

Nuamah:1986:PCS

[426] Nicholas N. N. N. Nuamah. Pooling cross section and time series data. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3): 345–351, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987750.

Rizvi:1986:MDM

[427] S. M. Rizvi. Manpower demand modelling. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (3):353–358, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987751.

Wheeler:1986:TCR

[428] Frederick P. Wheeler. Tracking changes in the rate of market growth. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):359–368, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987752.

Tarling:1986:SAC

[429] Roger Tarling. Statistical applications in criminology. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):369–388, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987753.

Anonymous:1986:BRB

[430] Anonymous. Book review: Applied Social Research Methods Series. Vol. 2: The Integrative Research Review; A Systematic Approach, by Harris M. Cooper; Applied Social Research Methods Series. Vol. 3: Methods for Policy Research, by Ann Majchrzak; Applied Social Research Methods Series. Vol. 4: Secondary Research; Information Sources and Methods, by David W. Stewart; Applied Social Research Methods Series. Vol. 5: Case Study Research: Design and Methods, by Robert K. Yin; Applied Social Research Methods Series. Vol. 6: Meta-Analytic Procedures for Social Research, by Robert Rosenthal. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):389–391, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987754.

Pridmore:1986:BRB

[431] W. A. Pridmore. Book review: Comparative Models for Electrical Load Forecasting, by D. W. Bunn; E. D. Farmer. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3): 391–392, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987755.

Bissell:1986:BRBb

[432] Derek Bissell. Book review: Statistics: A Spectator Sport, by Richard M. Jaeger. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):392, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987756.

Shoesmith:1986:BRBb

[433] Eddie Shoesmith. Book review: Bayesian Statistical Inference Quantitative Applications in the Social Sciences, Paper 43, by Gudmund R. Iverson. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (3):393, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987757.

Reese:1986:BRBc

[434] R. Allan Reese. Book review: Interpreting and Using Regression Quantitative Applications in the Social Sciences, Paper 29, by Christopher H. Achen; Linear Probability, Logit and Probit Models Quantitative Applications in the Social Sciences, Paper 45, by John H. Aldrich; Forrest D. Nelson. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):394, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987758.

Mellor:1986:BRB

[435] Chris Mellor. Book review: MINITAB Handbook, by Barbara F. Ryan; Brian L. Joiner; Thomas A. Ryan. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):395, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987759.

Sheldon:1986:BRBa

[436] Neil A. Sheldon. Book review: Naturalistic Enquiry, by Yvonna S. Lincoln; Egon G. Guba. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):395–396, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987760.

Davies:1986:BRB

[437] Neville Davies. Book review: Handbook of Statistics 5: Time Series in the Time Domain, by E. J. Hannan; P. R. Krishnaiah; M. M. Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):396–397, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987761.

Beaumont:1986:BRB

[438] Chris Beaumont. Book review: Exploring Data Tables, Trends and Shapes, by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):397, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987762.

Kemp:1986:BRBb

[439] C. D. Kemp. Book review: Statistical Inference Based on Ranks, by Thomas P. Hettmansperger. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):397–398, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987763.

Kemp:1986:BRBc

[440] A. W. Kemp. Book review: A Treatise on Generating Functions, by H. M. Srivastava; H. L. Manocha. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):398–399, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987764.

Chanter:1986:BRBb

[441] D. O. Chanter. Book review: Outliers in Statistical Data, by Vic Barnett; Toby Lewis. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):399, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987765.

Cooper:1986:BRB

[442] Philip J. Cooper. Book review: Linear Statistical Analysis of Discrete Data, by Mikel Aickin. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):400–401, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987766.

Higgins:1986:BRBa

[443] J. Higgins. Book review: Multicollinearity and Biased Estimation, by J. Gruber. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (3):401, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987767.

Shea:1986:BRB

[444] B. L. Shea. Book review: Time Series Package (TSPACK), by François S. Chaghaghi. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):401–402, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987768.

Smeeton:1986:BRB

[445] Nigel C. Smeeton. Book review: Exploring Clinical Methods for Social Research, by David N. Berg; Kenwyn K. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (3):402–403, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987769.

Higgins:1986:BRBb

[446] J. Higgins. Book review: Validity and the Research Process, by David Brunberg; Joseph E. McGrath. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):403–404, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987770.

Farebrother:1986:BRB

[447] R. W. Farebrother. Book review: Structural Sensitivity in Econometric Models, by Edwin Kuh; John W. Neese; Peter Hollinger. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):404, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987771.

Hawkins:1986:BRB

[448] Anne S. Hawkins. Book review: Microcomputer Methods for Social Scientists Quantitative Applications in the Social Sciences Series No. 40, by Philip A. Schrodt. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):404–405, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987772.

Sheldon:1986:BRBb

[449] Neil A. Sheldon. Book review: Advanced Mathematics 1, by C. W. Celia; A. T. F. Nice; K. F. Elliott. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):405, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987773.

Anonymous:1986:BMc

[450] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):??, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987774.

Anonymous:1986:FMd

[451] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):407–408, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987793.

OQuigley:1986:CSA

[452] John O’Quigley. The chi-squared approximation in the comparison of several proportions. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):409–412, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987794.

Oksanen:1986:NOS

[453] E. H. Oksanen. A note on observationspecific dummies and logit analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):

413–416, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987795.

Cox:1986:ARD

[454] C. Philip Cox and Richard A. Groeneveld. Analytic results on the difference between the G^{2} and χ^{2} test statistics in one degree of freedom cases. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4): 417–420, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987796.

Kinsella:1986:EMP

[455] A. Kinsella. Estimating method precisions. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (4):421–427, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987797.

Gibbs:1986:RTS

[456] M. Gibbs, R. K. Lindner, and A. J.

Fischer. Reliability of two survey techniques: A study of innovation discovery by farmers. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):429–439, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987798.

Romeu:1986:TES

[457] Jorge Luis Romeu. Teaching engineering statistics with simulation: A classroom experience. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):441–447, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987799.

Bora:1986:RCD

[458] A. C. Bora, M. Singh, and G. Singh. Robust change-over designs. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):449–454, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987800.

Mantel:1986:DEP

[459] Nathan Mantel and Samuel W. Greenhouse. On the Diaconis–Efron proposal for a volume test for contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (4):455–457, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987801.

Laschober:1986:AAA

[460] Paula J. Laschober and H. Tamura. Application of AID in analysis of energy survey data — a case study. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4): 459–469, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987802.

Shore:1986:ACF

[461] Haim Shore. Approximate closed form expressions for the decision variables of some tests related to the binomial distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):471–478, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987803.

Frisen:1986:UR

[462] M. Frisen. Unimodal regression. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):479–485, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987804.

Reese:1986:LEA

[463] Allan Reese. Letter to the Editor: An addendum. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):487, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987805.

Anonymous:1986:BMd

[464] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(4):??, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987806.

Anonymous:1986:FMe

[465] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):489–490, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987964.

Johnstone:1986:TST

[466] D. J. Johnstone, G. A. Barnard, and D. V. Lindley. Tests of significance in theory and practice. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):491–504, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987965.

Daffin:1986:MSA

[467] C. Daffin, P. Duncombe, Mary G. Hills, P. M. North, and G. B. Wetherill. A microcomputer survey analysis package designed for the developing countries. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5): 505–523, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987966.

Loukas:1986:CSG

[468] S. Loukas and C. D. Kemp. On the chi-square goodness-of-fit statistic for bivariate discrete distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):525–529, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987967.

Oshungade:1986:UPC

[469] Isaac Olayiwola Oshungade. Use of percentage change in small area statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5): 531–545, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987968.

Freeman:1986:PRA

[470] James M. Freeman. A piecewise regression analysis for firm-size distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (5):547–559, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987969.

Chaghaghi:1986:BRB

[471] François S. Chaghaghi. Book review:

Probability and Measure, by Patrick Billingsley. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):561, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987970.

Krotki:1986:BRBa

[472] Karol J. Krotki. Book review: Secondary Analysis of Survey Data Quantitiative Applications in the Social Sciences, Paper 53, by K. Jill Kielcolt; Laura E. Nathan. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):561–563, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987971.

Lawson:1986:BRB

[473] Andrew Lawson. Book review: Essential Statistics, by D. G. Rees. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):563, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987972.

Kinsella:1986:BRBb

[474] A. Kinsella. Book review: Finite Algorithms in Optimisation and Data Analysis, by M. R. Osborne. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):564, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987973.

Cheek:1986:BRB

[475] P. J. Cheek. Book review: Handbook of Statistics 3: Time Series in the Time Domain, by D. R. Brillinger; P. R. Krishnaiah. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):564–565, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987974.

Law:1986:BRB

[476] John Law. Book review: Robust Statistics — The Approach Based on Influence Functions, by F. R. Hampel; E. M. Ronchetti; P. J. Rousseuw; W. A. Stahel. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (5):565–566, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987975.

Preece:1986:BRB

[477] D. A. Preece. Book review: Multiple Regression in Hydrology, by R. L. Holder. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):566, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987976.

Gibson:1986:BRB

[478] Martin G. Gibson. Book review: Beyond ANOVA: Basics of Applied Statistics, by R. G. Miller. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):566–567, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987977.

Morgan:1986:BRBb

[479] D. D. V. Morgan. Book review: Handbook of Statistics 4: Nonparametric Methods, by P. R. Krishnaiah; P. K. Sen. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (5):567, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987978.

Horgan:1986:BRB

[480] Jane M. Horgan. Book review: Survey Design and Analysis, by F. R. Jolliffe. Journal of the Royal Statistical Society. Series D (The Statistician), 35 (5):568, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987979.

Hartzig:1986:BRB

[481] R. J. Hartzig. Book review: Statistical Analysis of Measurement Errors, by John L. Jaech. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):568–569, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987980.

Cuming:1986:BRB

[482] W. M. Cuming. Book review: Mastering Statistics with Your Microcomputer, by Conall Boyle. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):569–570, ???? 1986. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987981.

Kenett:1986:BRB

[483] Ron S. Kenett. Book review: Fundamentals of Queueing Theory, by Donald Gross; Carl M. Harris. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):570, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987982.

Krotki:1986:BRBb

[484] Karol J. Krotki. Book review: The Politics and Ethics of Fieldwork Qualitative Research Methods Series, Paper 3, by Maurice Punch. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):571, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987983.

Anonymous:1986:BMe

[485] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):572, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987984.

Anonymous:1986:VI

[486] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 35(5):573–578, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987963.

Anonymous:1987:FMa

[487] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):1–2, ???? 1987.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988266.

Robinson:1987:EUV

[488] Dorothy L. Robinson. Estimation and use of variance components. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):3–14, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988267.

Pereira:1987:EQU

[489] C. A. de B. Pereira and D. V. Lindley.

Examples questioning the use of partial likelihood. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):15–20, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988268.

Mantel:1987:TUP

[490] Nathan Mantel. Truncation as a unifying principle. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):21–24, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988269.

Prabhakaran:1987:VHC

[491] V. T. Prabhakaran and J. P. Jain.

On the variance of heritability coefficient obtained from intra-sire regression. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (1):25–26, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988270.

Katz:1987:PDP

[492] Leon Katz. Phi and the difference in proportions. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):27–29, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988271.

Dagpunar:1987:NMS

[493] J. S. Dagpunar. Nomograms for the manual sampling of random variates from gamma, normal and associated distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):31–36, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988272.

Dunsmore:1987:BAS

[494] Ian R. Dunsmore. A Bayesian alternative to the sign test in a matched pairs design. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):37–43, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988273.

Preece:1987:LSQ

[495] D. A. Preece. The language of size, quantity and comparison. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):45–54, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988274.

Hendl:1987:NET

[496] Jan Hendl. A note on exponential transformations. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):55–57, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988275.

Taylor:1987:BRB

[497] Eddie Taylor. Book review: ESRC Data Archive Catalogue, by Marcia Freed Taylor. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (1):59, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988276.

Rees:1987:BRBa

[498] D. G. Rees. Book review: Basic Statistics with Applications, by Edwin Mansfield. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):60, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988277.

Calder:1987:BRB

[499] Allan B. Calder. Book review: R. A.

Fisher: The Life of a Scientist, by R. A. Fisher; Joan Fisher Box. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):60–62, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988278.

Chadwick:1987:BRB

[500] Jeremy Chadwick. Book review: Meta-Analysis, by Frederic M. Wolf. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):62, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988279.

Garreau:1987:BRBa

[501] G. A. Garreau. Book review: Analyse Numérique des Équations Différentielles, by M. Crouzeix; A. L. Mignot. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):63, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988280.

Rees:1987:BRBb

[502] D. G. Rees. Book review: Linear Programming, An Introduction, by Bruce R. Feiring. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):63–64, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988281.

Kerdar:1987:BRB

[503] Nicholas Kerdar. Book review: The Best of Teaching Statistics, by Peter Holmes. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):64–65, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988282.

Barmby:1987:BRB

[504] Tim Barmby. Book review: Plain Figures, by Myra Chapman; Basil Mahon. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):65, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988283.

Kemp:1987:BRBa

[505] A. W. Kemp. Book review: Systems in Stochastic Equilibrium, by P. Whittle. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (1):65–66, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988284.

Kingman:1987:BRB

[506] J. F. C. Kingman. Book review: Brownian Motion and Stochastic Flow Systems, by J. Michael Harrison. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):66–67, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988285.

Hawkins:1987:BRB

[507] Anne S. Hawkins. Book review: Bandit Problems-Sequential Allocation of Experiments, by Donald A. Berry; Bert Fristedt. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):67, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988286.

Rowe:1987:BRB

[508] N. Rowe. Book review: Modern Analytical Techniques, by Frank Owen; Ron Jones. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):68, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988287.

Reese:1987:BRB

[509] Allan Reese. Book review: BTO Guide 22: Statististics for Ornithologists, by J. Fowler; L. Cohen. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):69, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988288.

Sparks:1987:BRBa

[510] T. Sparks. Book review: Understanding Regression Analysis: An Introductory Guide Quantitative Applications in the Social Sciences No. 57, by Larry D. Schroeder; David L. Sjoquist; Paula E. Stephan. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):70, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988289.

Elis-Williams:1987:BRB

[511] David Elis-Williams. Book review: The Quantum World, by J. C. Polkinghorne. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1): 70–71, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988290.

Anonymous:1987:BMa

[512] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):??, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988291.

Smith:1987:PNG

[513] A. F. M. Smith, A. M. Skene, J. E. H.

Shaw, and J. C. Naylor. Progress with numerical and graphical methods for practical Bayesian statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 75–82, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348499.

VanDijk:1987:ACP

[514] Herman K. Van Dijk, J. Peter Hop, and Adri S. Louter. An algorithm for the computation of posterior moments and densities using simple importance sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):83–90, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348500.

Dawid:1987:DAC

[515] A. P. Dawid. The difficulty about conjunction. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):91–97, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348501.

Evett:1987:BIF

[516] I. W. Evett. Bayesian inference and forensic science: Problems and perspectives. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):99–105, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348502.

Bermudez:1987:CPE

[517] J. D. Bermudez, J. M. Bernardo, and M. Sendra. Classification problems in education. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):107–113, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2348503.

Harrison:1987:PBF

[518] Jeff Harrison and Mike West. Practical Bayesian forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):115–125, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2348504.

Gilks:1987:SAH

[519] W. R. Gilks. Some applications of hierarchical models in kidney transplantation. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):127–136, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348505.

Bayarri:1987:BAS

[520] M. J. Bayarri and M. H. DeGroot. Bayesian analysis of selection models. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 137–146, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348506.

Shaw:1987:NBA

[521] J. E. H. Shaw. Numerical Bayesian analysis of some flexible regression models. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):147–153, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348507.

Adcock:1987:BAC

[522] C. J. Adcock. A Bayesian approach to calculating sample sizes for multinomial sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):155–159, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348508.

Lee:1987:AII

[523] T. D. Lee. Assessment of inter- and intra-laboratory variances: A Bayesian alternative to BS 5497. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):161–170, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348509.

VanDerMerwe:1987:BEB

[524] A. J. Van Der Merwe and P. C. N.

Groenewald. Bayes and empirical Bayes confidence intervals in applied research. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 171–179, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348510.

Berry:1987:SID

[525] Donald A. Berry. Statistical inference, designing clinical trials, and pharmaceutical company decisions. Journal of the Royal Statistical Society.

Series D (The Statistician), 36(2/3): 181–189, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348511.

Goldstein:1987:SAL

[526] Michael Goldstein. Systematic analysis of limited belief specifications. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 191–199, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348512.

Spiegelhalter:1987:CEP

[527] David J. Spiegelhalter. Coherent evidence propagation in expert systems. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 201–210, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348513.

Stewart:1987:HBA

[528] Leland Stewart. Hierarchical Bayesian analysis using Monte Carlo integration: Computing posterior distributions when there are many possible models. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 211–219, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348514.

Polasek:1987:BRE

[529] Wolfgang Polasek. Bounds on rounding errors in linear regression models. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):

221–227, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348515.

Achcar:1987:TSD

[530] J. A. Achcar, H. Bolfarine, and Luis R.

Pericchi. Transformation of survival data to an extreme value distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 229–234, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348516.

Hills:1987:RPI

[531] S. E. Hills. Reference priors and identifiability problems in non-linear models. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):235–240, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348517.

Naylor:1987:BAT

[532] J. C. Naylor. Bayesian alternatives to t-tests. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):241–246, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348518.

OHagan:1987:MCF

[533] A. O’Hagan. Monte Carlo is fundamentally unsound. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):247–249, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348519.

Makov:1987:BTM

[534] Udi E. Makov. A Bayesian treatment of the ‘missing suspect’ problem in forensic science. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):251–258, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348520.

Kestemont:1987:KDC

[535] Marie-Paule Kestemont. The Kolmogorov distance as comparison measure between parametric and nonparametric Bayesian predictions. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 259–264, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348521.

Marriott:1987:BNG

[536] John Marriott. Bayesian numerical and graphical methods for Box–Jenkins time series. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (2/3):265–268, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348522.

Gamerman:1987:ADS

[537] Dani Gamerman and Mike West. An application of dynamic survival models in unemployment studies. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 269–274, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348523.

Quintana:1987:AIE

[538] José M. Quintana and Mike West.

An analysis of international exchange rates using multivariate DLM’s. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3): 275–281, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348524.

Grieve:1987:ABS

[539] A. P. Grieve. Applications of Bayesian software: Two examples. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):283–288, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348525.

Mouchart:1987:BAL

[540] Michel Mouchart and Hugo Roche.

Bayesian analysis of load-curves through spline functions. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):289–296, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348526.

Anonymous:1987:BMb

[541] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):??, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348527.

Anonymous:1987:FMb

[542] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(2/3):??, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348498.

Hawkins:1987:SSC

[543] Douglas M. Hawkins. Self-starting Cusum charts for location and scale. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4): 299–316, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348827.

Conniffe:1987:EML

[544] Denis Conniffe. Expected maximum log likelihood estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):317–329, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348828.

Laurence:1987:IMP

[545] A. F. Laurence. Imputing molecular properties. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):331–343, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348829.

Kinsella:1987:EBA

[546] A. Kinsella. The ‘exact’ bootstrap approach to confidence intervals for the relative difference statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):345–347, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348830.

Bunday:1987:MLE

[547] Brian D. Bunday and Victor A. Kiri.

Maximum likelihood estimation — practical merits of variable metric optimisation methods. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):349–355, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348831.

Chatfield:1987:EMF

[548] C. Chatfield and M. G. Schimek. An example of model-formulation using IDA. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4): 357–363, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348832.

Caroni:1987:RIM

[549] C. Caroni. Residuals and influence in the multivariate linear model. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):365–370, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348833.

Genest:1987:GDA

[550] Christian Genest and Philip E. J. Green.

A graphical display of association in twoway contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):371–380, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348834.

Caudill:1987:DCM

[551] Steven B. Caudill. Dichotomous choice models and dummy variables. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):381–383, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348835.

Mehrotra:1987:UCS

[552] P. C. Mehrotra, A. K. Srivastava, and K. K. Tyagi. On unequal cluster sampling for fixed sample size. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):385–391, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348836.

Khalique:1987:CUF

[553] A. Khalique and I. H. Tajuddin. On the choice of using four moments or three moments and the left boundary for fitting a Pearsonian distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4): 393–395, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348837.

Preece:1987:GSP

[554] D. A. Preece. Good statistical practice. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):397–408, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348838.

Prabhakaran:1987:POP

[555] V. T. Prabhakaran and J. P. Jain. A procedure for obtaining the probability of negative estimates of heritability from sib analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):409–411, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348839.

Nuamah:1987:BRB

[556] Nicholas N. N. N. Nuamah. Book review: Understanding Statistics, by Arnold Naiman; Robert Rosenfeld; Gene Zirkel. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):413–414, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348840.

Peters:1987:BRB

[557] Bob Peters. Book review: Sequential Methods in Statistics, by G. Barrie Wetherill; Kevin D. Glazebrook. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):414, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348841.

Farebrother:1987:BRB

[558] R. W. Farebrother. Book review: The History of Statistics: The Measurement of Uncertainty Before 1900, by Stephen M. Stigler. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):414–415, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348842.

Jones:1987:BRB

[559] D. A. Jones. Book review: Empirical Processes with Applications to Statistics, by G. R. Shorack; J. A. Wellner. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):415–416, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348843.

Bell:1987:BRB

[560] J. F. Bell. Book review: Statistical Inference in Linear Models, by H. Burke; O. Burke. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):416, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2348844.

Kemp:1987:BRBb

[561] A. W. Kemp. Book review: Stochastic Modelling and Analysis: A Computational Approach, by H. C. Tijms. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):417, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348845.

Morgan:1987:BRB

[562] D. D. V. Morgan. Book review: Multivariate Analysis of Variance, by J. H. Bray; S. E. Maxwell. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):417–418, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348846.

Gibson:1987:BRB

[563] Martin Gibson. Book review: The Design and Analysis of Clinical Experiments, by J. L. Fleiss. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):418–419, ???? 1987.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348847.

Sparks:1987:BRBb

[564] T. Sparks. Book review: Probability and Random Variables, by G. P. Beaumont. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):419–420, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348848.

Kemp:1987:BRBc

[565] C. D. Kemp. Book review: Density Estimation for Statistics and Data Analysis, by B. W. Silverman. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):420–421, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348849.

Lynn:1987:BRB

[566] Peter Lynn. Book review: Monte Carlo Optimization, Simulation and Sensitivity of Queueing Networks, by Reuven Y. Rubinstein. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):421–422, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348850.

Aitken:1987:BRB

[567] C. G. G. Aitken. Book review: Statistics and the Law, by M. H. Degroot; S. E. Fienberg; J. E. Kadane. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):422–423, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348851.

Garreau:1987:BRBb

[568] G. A. Garreau. Book review: Real and Complex Analysis, by Walter Rudin. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):423, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348852.

Porter:1987:BRB

[569] M. A. Porter. Book review: Analysis of Experiments with Missing Data, by Yadolah Dodge. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):423–424, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348853.

Essi:1987:BRBa

[570] Nyong O. Essi. Book review: Real and Stochastic Analysis, by M. M. Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):424, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348854.

Stuart:1987:BRBa

[571] Michael Stuart. Book review: Reliability and Quality Control, by A. P. Basu. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):424–426, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348855.

Day:1987:BRB

[572] S. J. Day. Book review: Survey Questions: Handcrafting the Standardized Questionnaire. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):426, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348856.

Fry:1987:BRB

[573] T. R. L. Fry. Book review: The Statistical Analysis of Compositional Data, by J. Aitchison. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):427, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348857.

Sheldon:1987:BRB

[574] Neil A. Sheldon. Book review: Revision Statistics, by M. G. Godfrey; E. M. Roebuck; A. J. Sherlock. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):427–428, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348858.

Essi:1987:BRBb

[575] Nyong O. Essi. Book review: Evaluation: A Systematic Approach, by Peter H. Rossi; Howard E. Freeman. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4): 428–429, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348859.

Croucher:1987:BRB

[576] R. J. Croucher. Book review: Reliability and Life Testing, by S. K. Sinha. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):429, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348860.

Chaplin:1987:BRB

[577] J. Alan Chaplin. Book review: Advanced Calculus for Engineering and Science Student, by Ian S. Murphy. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4): 429–430, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348861.

Clarke:1987:BRB

[578] Michael D. Clarke. Book review: Stochastic Simulation, by Brian D. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):430–431, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348862.

Hart:1987:BRB

[579] Anna Hart. Book review: Cognition as Intuitive Statistics, by Gerd Gegerenzer; David J. Murray. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):431–432, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348863.

Law:1987:BRB

[580] John Law. Book review: Principal Component Analysis, by I. T. Jolliffe. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):432, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348864.

Kinsella:1987:BRB

[581] A. Kinsella. Book review: Modelling Financial Time Series, by Stephen Taylor. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):433, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348865.

Roweth:1987:BRB

[582] B. Roweth. Book review: New Tools for Social Scientists: Advances and Applications in Research Methods, by William D. Berry; Michael S. Lewis-Beck. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (4):433–434, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348866.

Stuart:1987:BRBb

[583] Michael Stuart. Book review: The Logic of Causal Order, by James A. Davis. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):435, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348867.

Harris:1987:BRB

[584] Ray Harris. Book review: The Vote Catcher, by Mike Carey. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):435–436, ???? 1987.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348868.

Anonymous:1987:BMc

[585] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):??, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348869.

Anonymous:1987:FMc

[586] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(4):??, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348826.

Anonymous:1987:FMd

[587] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):437–438, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2348655.

Morrison:1987:SE

[588] S. J. Morrison. SQC is not enough. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5): 439–464, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348656.

Baillie:1987:MAS

[589] David H. Baillie. Multivariate acceptance sampling — some applications to defence procurement. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):465–478, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348657. See

errata [635, 814].

Lobley:1987:IQS

[590] D. A. Lobley. Industry, quality and statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):479–485, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348658.

Julian:1987:SKI

[591] K. Julian. Statisticians — keep it simple. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (5):487–492, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348659.

Manns:1987:ASQ

[592] T. S. Manns and M. J. Coleman. An approach to software quality assurance training. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):493–498, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348660.

Willmott:1987:RSP

[593] Michael Willmott and Michael Warren. Reliability surveys and panels. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5): 499–511, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348661.

Alvey:1987:MCA

[594] Norman G. Alvey. The measurement of consumer acceptability. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):513–523, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348662.

Wankling:1987:TSA

[595] R. M. Wankling. Trading standards: An aid to quality. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):525–530, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348663.

Townshend:1987:SCR

[596] I. J. Townshend and A. F. Bissell. Sampling for clinical report auditing. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5): 531–539, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348664.

Korabinski:1987:TBS

[597] Athol A. Korabinski. Teaching basic statistical quality control to the shopfloor. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (5):541–553, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348665.

Caulcutt:1987:SIF

[598] Roland Caulcutt. Statistics in industry: A failure of communication. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):

555–560, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348666.

Neave:1987:DPM

[599] Henry R. Neave. Deming’s 14 points for management: Framework for success. Journal of the Royal Statistical Society. Series D (The Statistician), 36 (5):561–570, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348667.

Murphy:1987:SCV

[600] B. J. Murphy. Selecting out of control variables with the T^{2} multivariate quality control procedure. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):571–581, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348668.

Searstone:1987:OTP

[601] K. Searstone. The open tech project in quality assurance. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):583–587, ???? 1987. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348669.

Wetherill:1987:CAD

[602] G. Barrie Wetherill and M. E. Gerson.

Computer aids to data quality control. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5): 589–592, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348670.

Anonymous:1987:VI

[603] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):593–597, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2348654.

Anonymous:1987:BMd

[604] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 36(5):??, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348671.

Anonymous:1988:FMa

[605] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):1–2, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348372.

Lin:1988:ACC

[606] Jinn-Tyan Lin. Approximating the cumulative chi-square distribution and its inverse. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):3–5, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348373.

Jackson:1988:OLS

[607] John D. Jackson and James A. Dunlevy. Orthogonal least squares and the interchangeability of alternative proxy variables in the social sciences. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):7–14, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348374.

Butcher:1988:UPC

[608] R. Butcher. The use of the postcode address file as a sampling frame. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1): 15–24, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348375.

Moors:1988:QAK

[609] J. J. A. Moors. A quantile alternative for kurtosis. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):25–32, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348376.

Good:1988:MMS

[610] Frances Good. A method of moderation of school-based assessments: Some statistical considerations. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):33–49, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348377.

OQuigley:1988:GAP

[611] John O’Quigley and Claude Baudoin.

General approaches to the problem of bioequivalence. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):51–58, ???? 1988. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348378.

Mantel:1988:CHF

[612] Nathan Mantel and Donald M. Stablein. The crossing hazard function problem. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):59–64, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348379.

Caudill:1988:TEA

[613] Steven B. Caudill. Type I errors after preliminary tests for heteroscedasticity. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (1):65–68, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348380.

Morgan:1988:SFA

[614] D. D. V. Morgan and A. S. Monk. Sampling farm accidents. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):69–79, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348381.

Kemp:1988:BRBa

[615] A. W. Kemp. Book review: Applied Probability and Queues, by Soren Asmussen. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):81, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348382.

Shoesmith:1988:BRB

[616] Eddie Shoesmith. Book review: Empirical Model-Building and Response Surfaces, by G. E. P. Box; N. R. Draper.

Journal of the Royal Statistical Society. Series D (The Statistician), 37 (1):82, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348383.

Corbett:1988:BRB

[617] Margaret Corbett. Book review: Measurement Error Models, by Wayne A. Fuller. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):83, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348385.

Garreau:1988:BRB

[618] G. A. Garreau. Book review: The Problems of Mathematics, by Ian Stewart. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (1):83, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348384.

Farebrother:1988:BRB

[619] R. W. Farebrother. Book review: Specification Analysis in the Linear Model, by Maxwell L. King; David E. A. Giles. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):84, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348386.

Kemp:1988:BRBb

[620] C. D. Kemp. Book review: Spatial Variation, by Bertil Matérn. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):84–85, ???? 1988.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348387.

Marks:1988:BRB

[621] Peter Marks. Book review: Statistical Methods in Discrimination Litigation, by D. H. Kaye; Mikel Aickin. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):85, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348388.

Harris:1988:BRB

[622] Ray Harris. Book review: Directory of Statistical Microcomputer Software, by Wayne A. Woodward; Alan C. Elliott; Henry L. Gray. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):85–86, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2348389.

Kempson:1988:BRB

[623] Robert E. Kempson. Book review: A Genstat Primer, by A. J. Weekes. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1): 86–87, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348390.

Smeeton:1988:BRBa

[624] Nigel C. Smeeton. Book review: A Guide to Statistical Methods and to the Pertinent Literature, by Lothar Sachs. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (1):87, ???? 1988. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348391.

Rees:1988:BRBa

[625] D. G. Rees. Book review: Elements of Statistical Inference, by D. V. Huntsberger; P. P. Billingsley. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):88, ???? 1988. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348392.

Preece:1988:BRB

[626] D. A. Preece. Book review: Cyclic Designs, by J. A. John. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):88–89, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348393.

Bell:1988:BRBa

[627] J. F. Bell. Book review: Multivariate Statistical Simulations, by Mark E. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):89, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348394.

Sparks:1988:BRB

[628] T. Sparks. Book review: A Handbook of Introductory Statistical Methods, by C. Phillip Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):89–90, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348395.

Pigden:1988:BRB

[629] Norman Pigden. Book review: Multidimensional Scaling: History, Theory and Applications, by Forrest W. Young; Robert M. Hamer. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):90–92, ???? 1988. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348396.

Jefferson:1988:BRB

[630] Sheila Jefferson. Book review: Survey Research Designs: Towards a Better Understanding of Their Costs and Benefits, by R. W. Pearson; R. F. Boruch. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):92, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348397.

Smith:1988:BRB

[631] Graeme Smith. Book review: Statistics for Management, by Richard Levin. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1): 92–93, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348398.

Kenett:1988:BRB

[632] Ron S. Kenett. Book review: Statistics for Business and Economics, by Paul Newbold. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):93–94, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348399.

Basford:1988:BRB

[633] K. E. Basford. Book review: Design and Analysis of Experiments, by M. N. Das; N. C. Giri. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):94, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348400.

McConway:1988:BRB

[634] Kevin J. McConway. Book review: Decision Theory: An Introduction to the Mathematics of Rationality, by Simon French. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):94–95, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348401.

Baillie:1988:EMA

[635] David H. Baillie. Errata: Multivariate acceptance sampling — some applications to defence procurement. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):97, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348402. See

[589].

Kinsella:1988:EEB

[636] A. Kinsella. Errata: The ‘exact’ bootstrap approach to confidence intervals for the relative difference statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (1):97, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348403.

Anonymous:1988:BMa

[637] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(1):??, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348404.

Anonymous:1988:FMb

[638] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):99– 100, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348684.

Tong:1988:MSN

[639] Howell Tong and R. Moeanaddin. On multi-step non-linear least squares prediction. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (2):101–110, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2348685.

Newbold:1988:PPL

[640] Paul Newbold. Predictors projecting linear trends plus seasonal dummies. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2): 111–127, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348686.

Chatfield:1988:HWF

[641] Chris Chatfield and Mohammad Yar. Holt–Winters forecasting: Some practical issues. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):129–140, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348687.

Wu:1988:BPU

[642] Lilian Shiao-Yen Wu. Business planning under uncertainty: Quantifying variability. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2): 141–151, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348688.

Perriman:1988:IAD

[643] W. S. Perriman. Inference associated with differenced time series data on measurements with error on ash content of coal. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2): 153–164, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348689.

Horn:1988:PSF

[644] Paul S. Horn, Paul W. Britton, and Daniel F. Lewis. On the prediction of a single future observation from a possibly noisy sample. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):165–172, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348690.

Yeomans:1988:BFS

[645] K. A. Yeomans and M. T. Bendixen.

Business forecasting in south Africa: Practice and cases. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):173–186, December 1988.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348691.

Marriott:1988:ASA

[646] J. M. Marriott and A. R. Tremayne. Alternative statistical approaches to time series modelling for forecasting purposes. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (2):187–197, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2348692.

Davies:1988:API

[647] N. Davies, J. Pemberton, and J. D. Petruccelli. An automatic procedure for identification, estimation and forecasting univariate self exiting threshold autoregressive models. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):199– 204, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348693.

Adcock:1988:SSM

[648] C. J. Adcock. A service system model for sales forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):205–211, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348694.

Moeanaddin:1988:CLR

[649] R. Moeanaddin and Howell Tong. A comparison of likelihood ratio test and CUSUM test for threshold autoregression. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):

213–225, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348695.

Ameen:1988:SDE

[650] Jamal R. M. Ameen. Sequential discount estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):227–237, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348696.

Adcock:1988:ECB

[651] C. J. Adcock. Erratum: Corrections to a Bayesian approach to calculating sample sizes for multinomial sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2): 239, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348697.

Anonymous:1988:BMb

[652] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(2):240– 241, December 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348698.

Kemp:1988:RED

[653] C. David Kemp and Adrienne W. Kemp.

Rapid estimation for discrete distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (3):243–255, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348162.

Stevenson:1988:PAM

[654] W. G. Stevenson, J. R. Mallon, and F. Hepper. Practical aspects of monitoring equality of employment opportunity in a large organisation. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):257–270, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348163.

Agliari:1988:RIP

[655] A. Agliari and C. Calvi Parisetti. A-g reference informative prior: A note on Zellner’s g prior. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):271–275, ???? 1988. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348164.

Glasbey:1988:ERS

[656] C. A. Glasbey. Examples of regression with serially correlated errors. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):277–291, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2348165.

Herbert:1988:ENR

[657] John H. Herbert and Phillip S. Kott. An empirical note on regressions with and without a poorly measured variable using gas demand as a case study. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3): 293–298, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348166.

Rees:1988:CPV

[658] D. G. Rees and C. J. K. Henry. On comparing the predicted values from two simple linear regression lines. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):299–306, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348167.

Bacon-Shone:1988:BMG

[659] John Bacon-Shone. Bayesian methods for growth curves with capturerecapture. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):307–312, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348168.

Preece:1988:BHK

[660] D. A. Preece, G. J. S. Ross, and P. J. Kirby. Bortkewitsch’s horsekicks and the generalised linear model. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3): 313–318, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348169.

Wright:1988:EPS

[661] David E. Wright and Claire E. Hazelhurst. Estimation and prediction for a simple software reliability model. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):319–325, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348170.

Bellhouse:1988:VES

[662] D. R. Bellhouse and B. C. Sutradhar. Variance estimation for systematic sampling when autocorrelation is present. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3): 327–332, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348171.

VonCollani:1988:EPC

[663] Elart Von Collani. Economic product control by routine testing of small samples. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (3):333–341, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348172.

Giles:1988:BRBa

[664] Philip Giles. Book review: SuperCalc: Assignments and Solutions, by B. M. Garstang. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):343, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348173.

Lawson:1988:BRBa

[665] Andrew Lawson. Book review: Stochastic Analysis of Computer Storage, by O. I. Aven; E. G. Coffman,; Y. A. Kogan. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (3):343–344, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348174.

Colwell:1988:BRB

[666] D. J. Colwell and J. R. Gillett. Book review: Statistics Explained: Basic Concepts and Methods, by R. Kapadia; G. Anderson. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):344, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348175.

Smeeton:1988:BRBb

[667] Nigel C. Smeeton. Book review: The Spectral Analysis of Time Series, by I. G. Zurbenko. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):344–345, ???? 1988. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348176.

Giles:1988:BRBb

[668] Philip Giles. Book review: An Introduction to Data Handling in BBC Basic, by J. Gatenby. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):345, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348177.

Hawkins:1988:BRBa

[669] Anne S. Hawkins. Book review: Instructional Technology: Foundations, by R. M. Gagné. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):345–346, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348178.

North:1988:BRB

[670] Philip M. North. Book review: Introduction to Biostatistics, by R. R. Sokal; F. J. Rohlf. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):346–347, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348179.

Cohen:1988:BRB

[671] Geoff Cohen. Book review: Statistics Problems and Solutions, by E. E. Bassett; J. M. Bremner; I. T. Jolliffe; B. Jones; B. J. T. Morgan; P. M. North. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3): 347–348, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348180.

Waller:1988:BRB

[672] Paul Waller. Book review: Expert Systems and Artificial Intelligence in Decision Support Systems, by H. G. Sol; C. A. T. Takkenberg; P. F. De Vries Robbe. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (3):348–349, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348181.

Hawkins:1988:BRBb

[673] Anne Hawkins. Book review: Information Theory: Structural Models for Qualitative Data, by K. Krippendorff. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (3):349, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348182.

Dugard:1988:BRBa

[674] P. I. Dugard. Book review: Engineering Reliability and Risk in Water Resources, by L. Duckstein; E. J. Plate. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3): 349–350, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348183.

Lawson:1988:BRBb

[675] Andrew Lawson. Book review: Introduction to Statistical Inference, by J. C. Kiefer. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):350, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348184.

Smeeton:1988:BRBc

[676] Nigel C. Smeeton. Book review: Analysis of Variance, by G. R. Iverson; H. Norpoth. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):351–352, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348185.

Jeffers:1988:BRB

[677] J. N. R. Jeffers. Book review: Improving Study Skills, by I. Selmes. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):352, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348186.

Anonymous:1988:BMc

[678] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):??, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348187.

Anonymous:1988:FMc

[679] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(3):??, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348161.

Anonymous:1988:FMd

[680] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 353–356, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348758.

Kinsella:1988:TDD

[681] A. Kinsella and C. O’Boyle. Testing differential drug effects. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 357–361, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348759.

Paul:1988:MDA

[682] S. R. Paul and Nathan Mantel. Model dependent analyses of litter-depletion data. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (4/5):363–370, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348760.

Bertrand:1988:QMR

[683] Philip V. Bertrand and Roger L. Holder.

A quirk in multiple regression: The whole regression can be greater than the sum of its parts. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):371–374, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348761.

Schlittgen:1988:GFT

[684] Rainer Schlittgen. Goodness-of-fit tests for uniformity and normality based on the correlation coefficient. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 375–381, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348762.

Shih:1988:NPS

[685] Weichung J. Shih. Non-parametric simultaneous prediction intervals for quality control. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):383–386, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348763.

Fung:1988:RCI

[686] Karen Yuen Fung and Hilda Tam. Robust confidence intervals for comparing several treatment groups to a control group. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (4/5):387–399, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348764.

Fuchs:1988:ACS

[687] Camil Fuchs and Ron S. Kenett. Appraisal of ceramic substrates by multivariate tolerance regions. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):401–411, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348765.

Lin:1988:AHA

[688] Jinn-Tyan Lin. Alternatives to Hamakers’ approximations to the cumulative normal distribution and its inverse. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 413–414, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348766.

Cohen:1988:DAC

[689] L. Jonathan Cohen. The difficulty about conjunction in forensic proof. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 415–416, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348767.

McLachlan:1988:CSV

[690] G. J. McLachlan. On the choice of starting values for the EM algorithm in fitting mixture models. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):417–425, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348768.

Gart:1988:ACI

[691] John J. Gart and Jun-Mo Nam. Alternative confidence intervals for the relative difference. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):427–431, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348769.

Adcock:1988:BAC

[692] C. J. Adcock. A Bayesian approach to calculating sample sizes. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):433–439, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348770.

Conniffe:1988:OEM

[693] Denis Conniffe. Obtaining expected maximum log likelihood estimators. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 441–449, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348771.

Congdon:1988:DLW

[694] Peter Congdon. Deprivation in London wards: Mortality and unemployment trends in the 1980’s. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 451–472, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348772.

Moeanaddin:1988:ACC

[695] R. Moeanaddin and Howell Tong. Addendum & corrigendum: A comparison of likelihood ratio test and CUSUM test for threshold autoregression. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 473–474, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348773.

Lynn:1988:BRB

[696] Peter Lynn. Book review: Multiple Imputation for Nonresponse in Surveys, by D. B. Rubin. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):475–476, ???? 1988. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348774.

Porter:1988:BRB

[697] M. A. Porter. Book review: Graphical Exploratory Data Analysis, by S. H. C. Du Toit; A. G. W. Steyn; R. H. Stumpf. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):476–477, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348775.

Grove:1988:BRB

[698] Dan Grove. Book review: Akaike Information Criterion Statistics, by Y. Sakamoto; M. Ishiguro; G. Kitagawa. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 477–478, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348776.

Bell:1988:BRBb

[699] John F. Bell. Book review: Multilevel Models in Educational and Social Research, by H. Goldstein. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):478, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348777.

Woodward:1988:BRB

[700] Phil Woodward. Book review: Linear Models for Unbalanced Data, by S. R. Searle. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (4/5):478–479, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348778.

Ogilvie:1988:BRB

[701] Grant Ogilvie. Book review: Decision Science and Social Risk Management, by M. W. Merkhofer. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):480, ???? 1988. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348779.

Nuamah:1988:BRB

[702] N. N. N. N. Nuamah. Book review: Business and Economics Statistics with Computer Applications, by W. E. Becker; D. L. Harnett. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 480–481, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348780.

Darton:1988:BRB

[703] R. A. Darton. Book review: Statistical Design for Research, by L. Kish. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 481–483, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348781.

Cannings:1988:BRB

[704] C. Cannings. Book review: Search Problems, by R. Ahlswede; I. Wegener. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (4/5):483, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348782.

Rutherford:1988:BRB

[705] G. S. W. Rutherford. Book review: Multivariate Data Analysis with Readings, by J. F. Hair,; R. E. Anderson; R. L. Tatham. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):484, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348783.

Dunne:1988:BRB

[706] Peter C. Dunne. Book review: Stochastic Differential Systems, by V. S. Pugachev; I. N. Sinitsyn. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):484–485, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2348784.

Stuart:1988:BRB

[707] Michael Stuart. Book review: Design, Data and Analysis by Some Friends of Cuthbert Daniel, by C. L. Mallows. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 485–486, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348785.

Gibson:1988:BRB

[708] Martin G. Gibson. Book review: Model Selection, by H. Linhart; W. Zucchini. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 486–487, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348786.

Farrow:1988:BRB

[709] A. M. Farrow. Book review: Basic Operational Research, by P. G. Moore. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (4/5):487, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348787.

Anderson:1988:BRB

[710] Oliver D. Anderson. Book review:

Kendall’s Advanced Theory of Statistics, Volume 1: Distribution Theory, by A. Stuart; J. K. Ord. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):487–488, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348788.

MacArthur:1988:BRB

[711] Ewan W. MacArthur. Book review: GLIM: An Introduction, by M. J. R. Healy. Journal of the Royal Statistical Society. Series D (The Statistician), 37 (4/5):488–490, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348789.

Dugard:1988:BRBb

[712] P. I. Dugard. Book review: Compromise, Negotiation and Decision, by B. R. Munier; M. F. Shakun. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 490–491, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348790.

Speller:1988:BRB

[713] M. A. Speller. Book review: An Introduction to Linear Programming, by G. R. Walsh. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):491–492, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2348791.

Rees:1988:BRBb

[714] D. G. Rees. Book review: Elements of Statistics for the Life and Social Sciences, by B. M. Alfred. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 492–493, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348792.

Anonymous:1988:VI

[715] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5): 495–499, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348757.

Anonymous:1988:BMd

[716] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 37(4/5):??, ???? 1988. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348793.

Nayak:1989:EMC

[717] Tapan K. Nayak and Dayanand N. Naik.

Estimating multinomial cell probabilities under quadratic loss. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):3–10, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349013.

Donatos:1989:MCS

[718] George S. Donatos. A Monte Carlo study of k-class estimators for small samples with normal and non-normal disturbances. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):11–20, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2349014.

Behboodian:1989:NSS

[719] Javad Behboodian. A note on skewness and symmetry. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):21–23, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2349015.

Crombie:1989:EGP

[720] I. K. Crombie, W. C. S. Smith, J. M. Irving, and H. D. Tunstall-Pedoe. Experience with general practitioner lists as a sampling frame for a survey of cardiovascular risk factors. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):25–31, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349016.

Reese:1989:GEG

[721] R. Allan Reese. A graphics extension to the GLIM^{(}R)3 computer program. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (1):33–45, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349017.

Silver:1989:SFN

[722] M. S. Silver. Sampling frames and nonresponse problems in surveys of United Kingdom companies: A case study of the ESRC/DTI survey of management training. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):47–60, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349018.

Makuch:1989:GBM

[723] Robert W. Makuch, Michael A. Stephens, and Michael Escobar. Generalised binomial models to examine the historical control assumption in active control equivalence studies. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):61–70, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2349019.

Rodrigues:1989:SNP

[724] Josemar Rodrigues, Heleno Bolfarine, and José Galvão Leite. A simple nonparametric Bayes solution to the estimation of the size of a closed animal population. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):71–76, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349020.

Craggs:1989:BRB

[725] Carolyn Craggs. Book review: Statistics in Research: Basic Concepts and Techniques for Research Workers, by B. Ostle; L. C. Malone. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):77, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349021.

Iyer:1989:BRBa

[726] R. Iyer. Book review: Business Forecasting, by John E. Hanke; Arthur G. Reitsch. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):77, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349022.

Lynn:1989:BRBa

[727] Peter Lynn. Book review: Asymptotic Theory of Statistical Inference, by B. L. S. Prakasa Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):78, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2349023.

Rennolls:1989:BRB

[728] K. Rennolls. Book review: Genstat 5: An Introduction, by P. Lane; N. Galwey; N. Alvey. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):78–79, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349024.

Grey:1989:BRB

[729] D. R. Grey. Book review: Counterexamples in Probability, by Jordan Stoyanov. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (1):79–80, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349025.

Clarke:1989:BRBa

[730] Michael D. Clarke. Book review: Introductory Statistics, by Neil A. Weiss; Matthew J. Hassett. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):80, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349026.

Krotki:1989:BRBa

[731] Karol J. Krótki. Book review: Field Methods in Cross-Cultural Research, Cross-Cultural Research and Methodology Series, Volume 8, by Walter J. Lonner; John W. Berry. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):80–82, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349027.

Giddings:1989:BRB

[732] D. R. Giddings. Book review: Road Accident Statistics, by T. P. Hutchinson. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):82, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349028.

Gibson:1989:BRBa

[733] Martin G. Gibson. Book review: Statistical Analysis with Missing Data, by R. J. A. Little; D. B. Rubin. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):82–83, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349029.

Haugh:1989:BRBa

[734] Larry D. Haugh. Book review: Stochastic Hydrology, by Ian B. MacNeill; Gary J. Umphrey. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):83–84, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2349030.

Woods:1989:BRB

[735] A. J. Woods. Book review: Latent Class Analysis, by Allan L. Mc-Cutcheon. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):84–85, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349031.

Ruddock:1989:BRB

[736] Les Ruddock. Book review: New Directions for Database Systems, by Gad Ariav; James Clifford. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):85, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349032.

Rao:1989:BRB

[737] T. Subba Rao. Book review: Stochastic Methods in Structural Dynamics, by G. I. Schuëller; M. Sinozuka. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):85–86, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349033.

Shanbhag:1989:BRB

[738] D. N. Shanbhag. Book review: Characterization Problems Associated with the Exponential Distribution, by T. A. Azlarov; N. A. Volodin. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):86, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349034.

Conniffe:1989:COE

[739] Denis Conniffe. Corrigendum: Obtaining expected maximum log likelihood estimators. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):87, ???? 1989. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349035.

Anonymous:1989:BMa

[740] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):89–90, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2349036.

Anonymous:1989:FMa

[741] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(1):??, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349012.

Anonymous:1989:FMb

[742] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):91–92, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348305.

Wittkowski:1989:AUS

[743] Knut M. Wittkowski. An asymptotic UMP sign test for discretised data. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 93–96, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348306.

Gray:1989:URD

[744] J. Brian Gray. On the use of regression diagnostics. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):97–105, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348307.

Adcock:1989:PSP

[745] C. J. Adcock. Pre-sampling procedures. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 107–115, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348308.

Coe:1989:SSN

[746] R. D. Coe. The stochastic spreadsheet: A new statistical computing tool. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 117–120, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348309.

Paul:1989:MFA

[747] S. R. Paul and Nathan Mantel. Modelfree analyses of litter-depletion data. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 121–125, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348310.

Smeeton:1989:CDE

[748] Nigel C. Smeeton. Changes in the distribution of episodes of mental illness in a group of U.K. general practices. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 127–131, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348311.

Iyer:1989:BRBb

[749] Raja Iyer. Book review: Basic Forecasting Techniques, by David Johnson; Malcolm King. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):133, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348312.

Lee:1989:BRB

[750] P. M. Lee. Book review: Probability and Bayesian Statistics, by R. Viertl. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 133–134, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348313.

Peace:1989:BRB

[751] Ralph Peace. Book review: Understanding Data, by Peter Sprent. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):134, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348314.

Lynn:1989:BRBb

[752] Peter Lynn. Book review: Small Area Statistics: An International Symposium, by R. Platek; J. N. K. Rao; C. E. Sarndal; M. P. Singh. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):135, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348315.

Gibson:1989:BRBb

[753] Martin G. Gibson. Book review: How Many Subjects? Statistical Power Analysis in Research, by H. C. Kraeme; S. Thieman. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):135–136, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348316.

Krotki:1989:BRBb

[754] Karol J. Krotki. Book review: Systematic Fieldwork. Vol. I: Foundations of Ethnography and Interviewing, by Oswald Werner; G. Mark Schoepfle; Systematic Fieldwork. Vol. II: Ethnographic Analysis and Data Management, by Oswald Werner; G. Mark Schoepfle. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 136–137, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348317.

Harris:1989:BRB

[755] Deborah J. Harris. Book review: Multiple Comparisons, by Alan J. Klockers; Gilber Sax. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):137–138, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348318.

Laycock:1989:BRB

[756] P. J. Laycock. Book review: Robust Regression and Outlier Detection, by Peter J. Rousseuw; Annick M. Leroy. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):138, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348319.

Nuamah:1989:BRB

[757] Nicholas N. N. N. Nuamah. Book review: Sensitivity Analysis in Linear Regression, by Samprit Chatterjee; Ali S. Hadi. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (2):138–139, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348320.

Scanlan:1989:BRB

[758] Tony Scanlan. Book review: Elements of Statistical Computing: Numerical Computation, by Ronald A. Thisted. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 139–140, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348321.

Samson:1989:BRB

[759] W. B. Samson. Book review: Reliability Data Bases, by Aniello Amendola; Alfred Z. Keller. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):140, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348322.

Shoesmith:1989:BRBa

[760] Eddie Shoesmith. Book review: Statistical Reasoning, by Gary Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):140–141, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348323.

Bissell:1989:BRB

[761] Derek Bissell. Book review: Problem Solving — A Statistician’s Guide, by Christopher Chatfield. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):141–142, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348324.

Rees:1989:BRBa

[762] D. G. Rees. Book review: Introductory Medical Statistics, by R. F. Mould. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 142–143, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348325.

Mackenzie:1989:BRB

[763] I. G. Mackenzie. Book review: Review of Pharmacokinetics: Mathematical and Statistical Approaches to Metabolism and Distribution of Chemicals and Drugs, by A. Pecile; A. Resigno. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):143–144, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348326.

Reilly:1989:BRB

[764] James A. Reilly. Book review: Paradoxes in Probability Theory and Mathematical Statistics, by Gabor J. Szekely. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (2):144, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348327.

McAllister:1989:BRB

[765] S. A. McAllister. Book review: Automation in Survey Processing. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):144–145, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348328.

Cockram:1989:BRB

[766] Paul Cockram. Book review: Mathematics and Statistics for Accounting and Business Studies, by T. Bancroft; G. O’Sullivan. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):145–146, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348329.

Clarke:1989:BRBb

[767] G. M. Clarke. Book review: Introduction to Probability & Mathematical Statistics, by V. Fabian; J. Hannan. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):146–147, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348330.

Collins:1989:BRBa

[768] Alan J. Collins. Book review: Surface Organometallic Chemistry: Molecular Approaches to Surface Catalysis, by J.-M. Basset; B. C. Gates. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):147, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348331.

Clarke:1989:BRBc

[769] G. M. Clarke. Book review: W. G. Cochran’s Impact on Statistics, by P. S. R. S. Rao; J. Sedransk. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):148, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348332.

Rees:1989:BRBb

[770] D. G. Rees. Book review: Statistics for Health Management and Research, by Mark Woodward; Lesley M. A. Francis. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (2):148–149, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348333.

Lynn:1989:BRBc

[771] Peter Lynn. Book review: New Developments in Statistics for Psychology and the Social Sciences, by A. D. Lovie. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 149–150, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348334.

Beaumont:1989:BRB

[772] Chris Beaumont. Book review: External Trade Indices. Eurostat News Special Edition. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):150, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348335.

Kemp:1989:BRB

[773] Adrienne W. Kemp. Book review:

Means and Their Inequalities, by P. S. Bullen; D. S. Mitrinović; P. M. Vasić. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2): 150–151, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348336.

Anonymous:1989:BMb

[774] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(2):??, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348337.

Logothetis:1989:ENP

[775] N. Logothetis. Establishing a noise performance measure. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):155–174, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348871.

Kinsella:1989:BBM

[776] A. Kinsella. Bootstrapping a bioequivalence measure. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):175–179, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348872.

Boland:1989:MSC

[777] Philip J. Boland. Majority systems and the Condorcet jury theorem. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):181–189, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348873.

Ahmed:1989:PSE

[778] Ejaz Ahmed. On the probability of selecting extreme value populations associated with the smallest location parameters. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (3):191–195, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348874.

Ipinyomi:1989:NRC

[779] R. A. Ipinyomi and B. L. Adeleke. Nested row-column designs. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):197–202, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2348875.

Caudill:1989:MME

[780] Steven B. Caudill and John D. Jackson. Measuring marginal effects in limited dependent variable models. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):203–206, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348876.

Corbett:1989:BRB

[781] Margaret Corbett. Book review: Statistics & Data Analysis: An Introduction, by Andrew F. Siegel. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):207, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348877.

Shoesmith:1989:BRBb

[782] Eddie Shoesmith. Book review: Exploring Statistics with the IBM(R) PC, by David P. Doane. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):207–209, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348878.

Clifford:1989:BRB

[783] Frank Clifford. Book review: Challenges and Responses in Mathematics, by Geoffrey Howson. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):209, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2348879.

Basford:1989:BRB

[784] K. E. Basford. Book review: Comparative Experiments with Field Crops, by G. V. Dyke. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):209–210, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348880.

Bell:1989:BRB

[785] J. F. Bell. Book review: Plane Answers to Complex Questions: The Theory of Linear Models, by Ronald Christensen. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):210, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2348881.

Haugh:1989:BRBb

[786] Larry D. Haugh. Book review: Tables for Selecting the Best Population, by K. L. Butler; D. G. Butler. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):211, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2348882.

Metcalfe:1989:BRB

[787] A. V. Metcalfe. Book review: Tensor Methods in Statistics, by P. McCullagh. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (3):211–212, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348883.

Collins:1989:BRBb

[788] Alan J. Collins. Book review: Interfaces Under Laser Irradiation, by L. D. Laude; D. Bauerle; M. Wautelet. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (3):212, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348884.

Porter:1989:BRB

[789] Max Porter. Book review: Characterization and Analysis of Block Designs, by A. K. Nigam; P. D. Puri; V. K. Gupta. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (3):212–213, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2348885.

Prys-Williams:1989:BRB

[790] Allan G. Prys-Williams. Book review:

Membership Roles in Fields Research, by Patricia A. Adler; Peter Adler. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):213, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348886.

Spurr:1989:BRB

[791] B. D. Spurr. Book review: The Inverse Gaussian Distribution, by Raj S. Chhikara; J. Leroy Folks. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):213–214, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348887.

Anonymous:1989:BMc

[792] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):??, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348888.

Anonymous:1989:FMc

[793] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(3):??, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2348870.

Anonymous:1989:FMd

[794] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):215–216, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2349053.

Ng:1989:EPM

[795] Vee Ming Ng. On an estimation problem in multiple inspections. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):217–219, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349054.

Woodward:1989:CBM

[796] Mark Woodward. A computer-based method for determining optimal sample size with medical applications. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):221–226, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2349055.

VanBatenburg:1989:BDS

[797] Paul C. Van Batenburg and J. Kriens. Bayesian discovery sampling: A simple model of Bayesian inference in auditing. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (4):227–233, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349056.

Wurst:1989:CSS

[798] John Wurst, John Neter, and James Godfrey. Comparison of sieve sampling with random and cell sampling of monetary units. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):235–249, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349057.

Osanaiye:1989:SNM

[799] P. A. Osanaiye and C. O. Talabi. On some non-manufacturing applications of counted data cumulative sum (CUSUM) control chart schemes. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):251–257, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349058.

Pettitt:1989:RRA

[800] A. N. Pettitt and I. Bin Daud. Roughand-ready approximations for diagnostic statistics for proportional hazards regression. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):259–274, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349059.

Metcalfe:1989:REU

[801] A. V. Metcalfe. Ratio estimators with unequal probabilities sampling: A practical application. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):275–280, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349060.

Oshungade:1989:SMH

[802] Isaac Olayiwola Oshungade. Some methods of handling item non-response in categorical data. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):281–296, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349061.

Silver:1989:CPI

[803] Mick Silver. Consumer price indices and parallel markets. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):297–304, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349062.

Sutton:1989:BRB

[804] Chris Sutton. Book review: Statistics for Biologists, by R. C. Campbell. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (4):305, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349063.

Krotki:1989:BRBc

[805] Karol J. Krótki. Book review: Population Statistics: A Review of UK Sources, by Bernard Benjamin. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):305–307, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349064.

Clarke:1989:BRBd

[806] G. M. Clarke. Book review: The Method of Paired Comparisons, by H. A. David. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4): 307–308, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349065.

Gray:1989:BRB

[807] C. T. Gray. Book review: Applications of Fibonacci Numbers, by A.

N. Philippou; A. F. Horadam; G. E. Bergum. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):308, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349066.

Selby:1989:BRB

[808] Michael J. P. Selby. Book review:

Stochastic Equations for Complex Systems, by A. V. Skorohod. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):308–309, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349067.

Iyer:1989:BRBc

[809] R. Iyer. Book review: New Statistical Procedures for the Social Sciences, by Rand R. Wilcox. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):309–310, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349068.

Marriott:1989:BRB

[810] Alan Marriott. Book review: The Problem of The Reign of King Edward III: A Statistical Approach, by Eliot Slater. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):310, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2349069.

Smith:1989:BRB

[811] M. E. Smith and I. G. Mackenzie. Book review: Statistics with Vague Data, by Rudolf Kruse; Klaus Dieter Meyer.

Journal of the Royal Statistical Society. Series D (The Statistician), 38(4): 310–311, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349070.

Clarke:1989:BRBe

[812] G. M. Clarke. Book review: Nonnegative Matrices and Applicable Topics in Linear Algebra, by A. Graham. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4): 311–312, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349071.

Krotki:1989:BRBd

[813] Karol J. Krótki. Book review: The American Census: A Social History, by Margo J. Anderson. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):312–314, ???? 1989. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2349072.

Baillie:1989:CMA

[814] David H. Baillie. Corrigendum: Multivariate acceptance sampling — some applications to defence procurement. Journal of the Royal Statistical Society. Series D (The Statistician), 38 (4):315, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2349073. See [589].

Anonymous:1989:VI

[815] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):317–320, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2349052.

Anonymous:1989:BMd

[816] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):??, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2349074.